### v1

# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

Abstract
Motivated by the compelling need for coherent atom transport in a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L. Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, triangular, honeycomb, dimerized, and checkerboard lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms. At the outset, we propose the appropriate arrangements of acousto-optic modulators that give rise to a frequency imbalance between counterpropagating laser beams, thus leading to the dynamical-lattice effect in an arbitrary direction in the lattice plane. We subsequently obtain the dynamical-lattice trajectories that enable atom transport using shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type. We then quantify the resulting atom dynamics using the transport fidelity computed from the numerical solutions of the relevant time-dependent Schrödinger equation. We do so for various choices of the system parameters and transport directions, finding favorable results for the achievable transport times and robustness of the resulting transport to various experimental imperfections.
I Introduction
Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic ; Ton . As a result, coherent atom transport Qic enabled by moving the confining optical trap – either an optical lattice, an array of optical tweezers, or their combination – has lately been attracting considerable attention, on the experimental- and theoretical sides alike. While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STAf ; ContRev .
Coherent single-atom transport in an optically-trapped system entails a final atomic state that approximates - as closely as possible - the initial state in the comoving reference frame (i.e. the rest frame of the moving optical trap), up to an irrelevant global phase. This last requirement, tantamount to demanding minimization - ideally a complete absence - of vibrational excitations at the final point of transport, does not, however, necessarily rule out transient excitations at intermediate times Tor . It is this last aspect that motivates one to design moving-trap trajectories using STA-based control protocols STA ; STAf . Generally speaking, such protocols typically lead to the same final states as their adiabatic counterparts, but often require significantly shorter times, thus alleviating the debilitating effects of noise and decoherence. Owing to the fact that adiabatic control-parameter variations typically render some dynamical properties of the system invariant, inverse-engineering schemes based on Lewis-Riesenfeld invariants (LRIs) Lew (a) constitute the most pervasive subclass of STA quantum-control protocols.
The early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework. However, given that such simplifications typically do not apply in realistic systems - where there is either a significant coupling between the longitudinal and transverse degrees of freedom, or the relevant confining potential is anharmonic Zha subsequent theoretical studies have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hau and double-well  HauSto types. Atom transport in such systems is enabled by the dynamical-lattice effect, which is induced by appropriately arranged pairs of acousto-optic modulators (AOMs) that give rise to a frequency imbalance between two counterpropagating laser beams. These studies of atom transport in higher-dimensional optical lattices Sto (b); Hof were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials. The results obtained using this method are comparable to those corresponding to the quantum-speed limit HauSto .
In this paper, we aim to demonstrate – using the STA approach – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family of lattices includes square, triangular, dimerized, honeycomb, and checkerboard lattices as its special cases. Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie ; these gates originate from exchange interactions between qubits Heu ; Heu (2); Sto (e); Hay .
We start by proposing specific arrangements of AOMs that enable the dynamical-lattice effect in an arbitrary direction in the relevant lattice plane. We then set out to model the ensuing coherent atom transport by employing STA-type protocols, here based on inverse-engineering techniques involving LRIs. Having obtained the STA dynamical-lattice trajectories, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM). We then quantify the efficiency of STA-based atomic transport through the corresponding transport fidelity for a broad range of values of the relevant system parameters.
In addition to finding STA solutions to the atom-transport problem at hand, we devote significant attention to analyzing their sensitivity to errors, i.e. their robustness to various experimental imperfections, such as deviations in the control amplitudes, the duration of control, the unavoidable presence of noise, etc.
The remainder of this paper is organized as follows. We start Sec. II by providing the basic background of the atom-transport problem at hand, including the Hamiltonian describing such transport and the essential properties of the relevant family of optical lattices; we then specify concrete arrangements of AOMs that enable atom transport in the system under consideration. Section III is devoted to a short survey of the general STA-related results of relevance for the problem at hand. In the following, Sec. IV, we describe how these general results are utilized for obtaining STA dynamical-lattice trajectories and analyze typical examples thereof. Section V is devoted to the description of our approach for the treatment of atomic dynamics. In Sec. VI, we discuss the obtained results for the atom-transport fidelities. To end with, a short summary of the paper - along with the main conclusions drawn - is provided in Sec. VII. In Appendix A we briefly review the application of the FSOM for solving TDSEs, while Appendix B contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems.
II System and atom transport
To set the stage for further discussion, we introduce the optical-lattice system to be considered in what follows and the principal objectives of our investigation. In particular, in Sec. II.1 below we briefly formulate the problem of atom transport and the characteristic spatial, temporal, and energy scales. We then describe the main features and parameters characterizing the relevant family of 2D optical lattices (Sec. II.2), followed by a discussion of specific configurations of AOMs that give rise to the dynamical-lattice effect in this system in different directions in the xyx-y plane (Sec. II.3).
II.1 Atom transport in optical lattices
The problem to be discussed in what follows is that of transporting an atom of mass $m$ (e.g. of 40K), located at the time $t=0$ in one of the minima of a 2D optical-lattice potential, to another one of its minima, which is situated at a distant location in the xyx-y plane. In keeping with the conventional atom-transport phenomenology, we assume hereafter that the transport distance $d$ is at least an order of magnitude larger than the linear dimension (i.e. the spatial extent) $l_{0}$ of the atomic ground-state wave packet, that is, $d\gtrsim 10\>l_{0}$. The principal assumption inherent to atom transport in optical lattices is that an atom initially resides in the ground state of the potential well corresponding to a lattice-potential minimum and – having covered the distance $d$ at the time $t=t_{f}$ – also finds itself in the ground state of the potential well at its destination. In other words, the atom transport process amounts ideally to a perfect state transfer – from one ground state to its spatially dislocated counterpart described by the same wave function (up to a physically unobservable global phase) – and is not being accompanied by the creation of any excitations.
It will be assumed in this work that atom transport occurs by way of the dynamical-lattice effect; in the simplest scenario this effect is enabled by the interference of two counterpropagating laser beams with a mutual frequency imbalance. Such an imbalance can straightforwardly be created through the use of appropriately arranged pairs of AOMs (see Sec. II.3 below).
The generic form of a time-dependent Hamiltonian describing single-atom transport in the xyx-y plane of a 2D optical lattice with the potential $U_{\mathrm{L}}(\mathbf{r})\equiv U_{\mathrm{L}}(x,y)$ reads
$H_{\textrm{L}}(t)=--\frac{\hbar^{2}}{2m}\>\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+U_{\mathrm{L}}\left[\mathbf{r}-\mathbf{q}_{0}(t)\right]\>,$ (1)
where $\mathbf{q}_{0}(t)\equiv\left[q_{0}^{\mathrm{x}}(t),q_{0}^{\mathrm{y}}(t)\right]^{\intercal}$ is the time-dependent 2D vector describing the dynamical-lattice trajectory, i.e. the path of a potential minimum. The TDSE corresponding to this last Hamiltonian is typically discussed in the comoving frame. The wave function $\Phi(\mathbf{r},t)$ of an atom in this last reference frame is obtained from its lab-frame counterpart $\Psi(\mathbf{r},t)$ via a generalized Galilean transformation Got . More precisely, $\Phi(\mathbf{r},t)\equiv\hat{\mathcal{U}}(t)\Psi(\mathbf{r},t)$, where in the Hilbert space of the system at hand the unitary operator $\hat{\mathcal{U}}(t)$ is given by (see, e.g. SLK )
$\hat{\mathcal{U}}(t)=e^{i\hat{\mathbf{p}}\cdot\mathbf{q}_{0}(t)/\hbar}\>e^{-im\hat{\mathbf{r}}\cdot\dot{\mathbf{q}}_{0}(t)/\hbar}\>,$ (2)
with $\hat{\mathbf{r}}$ and $\hat{\mathbf{p}}$ being the single-atom (2D) position and momentum operators, respectively.
Based on the Hamiltonian in Eq. (1), in Sec. IV below we discuss how – using an STA-based method – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration. In particular, both of our approaches (see Sec. II.3 below) enable atom transport along an arbitrary direction in the xyx-y plane; by way of example, we will mostly be interested in transport along the $x$-, $y$-, and diagonal directions, with the respective transport distances being denoted by $d_{\mathrm{x}}$, $d_{\mathrm{y}}$, and $d_{\mathrm{r}}$.
Before closing these preliminary considerations, it is prudent to specify the characteristic time-, length-, and energy scales in the problem under consideration. In particular, the oscillation periods $T_{\mathrm{x}}\equiv 2\pi/\omega_{\mathrm{x}}$ and $T_{\mathrm{y}}\equiv 2\pi/\omega_{\mathrm{y}}$, corresponding to the harmonic approximation of the full optical potential [cf. Eq. (7) in Sec. II.2 below] in the $x$- and $y$ directions, will serve as the relevant time scales of atom transport in the respective directions. Likewise, the harmonic-oscillator zero-point lengths $l_{\textrm{x}}\equiv\sqrt{\hbar/(2m\omega_{\mathrm{x}})}$ and $l_{\textrm{y}}\equiv\sqrt{\hbar/(2m\omega_{\mathrm{y}})}$ will play the role of the characteristic length scales. Finally, the recoil energy $E_{\textrm{R}}\equiv\hbar^{2}k_{\textrm{L}}^{2}/(2m)$, where $k_{\textrm{L}}$ is the magnitude of the relevant laser wave vector $\mathbf{k}_{\textrm{L}}$, will serve as the characteristic energy scale.
II.2 Adjustable family of 2D optical lattices
In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remainder of this work. In order to provide a more concise and clearer explanation of the envisioned setup, we hereafter restrict our discussion to the strong confining potentials in the xyx-y plane; we neglect the minor modulations arising from the ever-present confinement of the system in the $z$ direction, which in the anticipated physical realization is close to being harmonic.
The sought-after adjustable family of 2D optical lattices is constructed from an arrangement of three distinct retro-reflected laser beams of linear polarization and wavelength $\lambda_{\mathrm{L}}$ Tarr ; therefore, the entire system can be parametrized by a single wave number $k_{\mathrm{L}}\equiv 2\pi/\lambda_{\mathrm{L}}$. These three beams will be denoted by $\overline{X}$, $X$, and $Y$ in the following (for an illustration, see Fig. 1), with $I_{\overline{X}}$, $I_{X}$, and $I_{Y}$ being the respective beam intensities.
In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis (cf. Fig. 1). In particular, the interference of two perpendicular beams – namely, beams $X$ and $Y$ – creates a checkerboard lattice with period $\lambda_{\mathrm{L}}/\sqrt{2}$. At the same time, a third beam, $\overline{X}$ – which is collinear with $X$, but detuned from it by a frequency $\delta$ – gives rise to an additional standing wave with a period of $\lambda_{\mathrm{L}}/2$.
The potential that corresponds to the envisioned adjustable family of 2D optical lattices can be derived starting from the general expression for the light intensity $I(x,y)$ in terms of the electric field $\mathbf{E}(x,y)$ Sho
$I(x,y)=c\epsilon_{0}\>|\mathbf{E}(x,y)|^{2}\>.$ (3)
By making use of this last expression, one finds that the intensity pattern corresponding to the checkerboard lattice is given by
$I_{1}(x,y)=$ - $I_{X}\,\sin^{2}(k_{\mathrm{L}}x)-I_{Y}\,\cos^{2}(k_{\mathrm{L}}y)$
- $2\alpha\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $\phi$ corresponds to the relative phase between the two arms, while $\alpha$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting. The possible sources of experimental imperfections are, for instance, imperfect coherence and mode-mismatching. In what follows, we will restrict our analysis to high-contrast regimes of $\alpha=0.9$, which is readily achievable in experimental realizations Tarr .
Figure 1: Schematic of a setup that enables the adjustable optical-lattice formed from three folded retro-reflected beams. The direct digital synthesizer (DDS) is used to control the frequency change induced by the acousto-optic modulator (AOM), giving rise to the phase $\theta$. The half-wave plates (HWP) can be adjusted in their angular orientation to set the specific intensities wanted for the three laser beams. The red dot indicates the center of the optical-lattice potential.
By once again employing the general expression for laser-beam intensity [cf. Eq. (3)], one finds that the laser-intensity pattern corresponding to beam $\overline{X}$ is given by
$I_{2}(x,y)=-I_{\overline{X}}\,\cos^{2}(k_{\mathrm{L}}x+\theta/2)\>,$ (5)
where the phase $\theta$ determines the relative position of the checkerboard pattern and the 1D standing wave; physically, $\theta$ originates from a detuning between the two laser beams along the $x$ axis, i.e. beams $\overline{X}$ and $X$ (for a schematic illustration, see Fig. 1 below).
From Eqs. (II.2) and (5) it is straightforward to obtain the expression for the total laser-beam intensity $I_{\mathrm{tot}}(x,y)=I_{1}(x,y)+I_{2}(x,y)$. It then immediately follows that the corresponding total optical-lattice potential $U_{\mathrm{L}}(x,y)\>\propto\>I_{\mathrm{tot}}(x,y)$, describing the envisioned family of 2D lattices, in its most general form reads
$U_{\mathrm{L}}(x,y)=$ - $U_{X}\,\sin^{2}(k_{\mathrm{L}}x)-U_{Y}\,\cos^{2}(k_{\mathrm{L}}y)$
- $U_{\overline{X}}\,\cos^{2}(k_{\mathrm{L}}x+\theta/2)$
- $2\alpha\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ are the three single-beam lattice depths (proportional to the laser-beam intensities). This last optical-lattice potential is closely related to the potential utilized in the experimental realization of an adjustable family of 2D optical-lattice structures by Tarruell et al. Tarr . For the sake of simplifying the analysis of coherent atom transport in the present work, we make the following two changes compared to Tarr . Firstly, we displace the $x$ coordinate by the distance $\pi/(2k_{\mathrm{L}})$ [i.e. make the replacement $x\rightarrow x-\pi/(2k_{\mathrm{L}})$]. Secondly, we adopt a somewhat different convention for the phase $\theta$; more specifically yet, we make the replacement $\theta\rightarrow\theta+\pi$.
The general optical-lattice potential in Eq. (II.2) gives rise – by adjusting the relative magnitude of the potential depths $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ – to several different lattice configurations; their corresponding potentials – as well as their respective periodically-repeating units (i.e. unit cells) – are visualized in Fig. 2. Among them, the following four lattice structures can be realized: square, dimerized, honeycomb, and 1D chains (cf. Fig. 2).
Figure 2: Density plots of the optical potential for the (a) square, (b) dimerized, (c) honeycomb, and (d) 1D-chains, optical lattices, obtainable by adjusting the different lattice depths present in the expression for $U_{\mathrm{L}}(x,y)$ [cf. Eq. (II.2)]. For each lattice structure, the corresponding unit cell is indicated by the dashed white line. The point $(x_{0},y_{0})=(0,0)$, around which the harmonic approximation of the full potential is developed, lies in the center od the unit cell.
Having analyzed the full, anharmonic optical-lattice potential of the system at hand [cf. Eq. (II.2)], it is pertinent to find at this point its harmonic counterpart, i.e. its approximate form, with terms up to quadratic order in $x$ and $y$. This harmonic potential, to be denoted by $V_{\mathrm{L}}(x,y)$, is of utmost importance for our further discussion of atom transport (see Secs. III.2 and IV below).
Given that the adjustable optical-lattice potential in Eq. (II.2) has minima around $(x_{0},y_{0})=(0,0)$ for all previously discussed configurations (cf. Fig. 2), we will approximate the full potential around the origin. We do so assuming that $k_{\mathrm{L}}|x|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$. Based on these assumptions, we straightforwardly obtain
$V_{\mathrm{L}}=-V_{\mathrm{d},0}+ma_{\mathrm{x}}x+\frac{m}{2}\left(\omega_{\mathrm{x}}^{2}x^{2}+\omega_{\mathrm{y}}^{2}y^{2}\right)\>,$ (7)
where $V_{\mathrm{d},0}$ depends on $\theta$ and is given by
$V_{\mathrm{d},0}=U_{Y}+U_{\overline{X}}\cos^{2}\left(\frac{\theta}{2}\right)\>.$ (8)
The squared frequencies in Eq. (7) are given by
$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\>,$ (9)
$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\>,$
while the expression for the factor $a_{\mathrm{x}}$, which has dimensions of acceleration, reads
$a_{\mathrm{x}}=\frac{U_{\overline{X}}k_{\mathrm{L}}}{m}\left(\sin\theta-2\alpha\>\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}}\cos\phi\right)\>.$ (10)
Therefore, the approximated potential has a term linear in $x$. While the expansions of the conventional optical-lattice potentials do not contain such terms, these unusual terms – linear in the spatial coordinate – are also present in the case of the double-well optical lattice HauSto .
Another important property of the approximated potential $V_{\mathrm{L}}$ in Eq. (7), of interest for our further discussion, is that – by contrast to the total lattice potential, which is nonseparable in the xyx-y plane (i.e. it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by this approximated potential.
II.3 AOM configurations for atom transport
To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$, $y$, and diagonal ($r$) directions.
To be able to appreciate the specific configurations of AOMs that give rise to the dynamical-lattice effect in the system at hand (see Fig. 3 below), it is helpful to first recall how the use of an AOM (also known as a Bragg cell) gives rise to the dynamical-lattice effect. AOMs lead to a shift of an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal. Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to Bragg diffraction. Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua . In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs. While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wave-numbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively. This imbalance engenders a dynamical (moving) optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wave number HauSto .
Figure 3: Adapted schematic of a setup that enables atom transport for the adjustable optical-lattice. The direct digital synthesizer (DDS) is used to control the frequency change induced by the acousto-optic modulator (AOM), giving rise to the dynamical-lattice velocity $v$. The red dot indicates the center of the optical-lattice potential.
Having described the general procedure for incorporating AOMs into the system to enable atom transport, we now discuss how this approach can be implemented in the setup for the adjustable potential shown in Fig. 1. Currently, a single AOM is employed to tune the phase shift $\theta$ present in the optical-lattice potential of Eq. II.2. However, it is important to clarify that this particular AOM does not participate in the shifting procedure. To enable controlled lattice translation, each arm of the setup must first be modified so that the beams are not simply reflected by a mirror at the end of each arm. Instead, additional beam splitters and mirrors must be incorporated to separate the incoming and reflected components in each arm, allowing them to be treated independently. An illustration of the modified configuration for the relevant region is provided in Fig. 3. Once the four incoming beams are separated, individual AOMs can be placed in each path to control the frequency of the corresponding beam. In this way, a dynamical optical lattice can be generated in each arm through an induced frequency difference $\Delta f$. The resulting transport directions can then be represented as a superposition of displacements along each axis. For ideal diagonal transport, the frequency differences in both arms must be tuned identically.
III STA: survey of results relevant for atom transport
To set the stage for modeling coherent atom transport in the family of optical lattices under consideration using the STA- and eSTA methods (see Sec. IV below), we first provide a short survey of these methods, with emphasis on results of relevance for atom transport. We start with a short introduction into the subclass of STA methods that employ inverse engineering based on LRIs (Sec. III.1) and then specialize our discussion to the problem of harmonic transport (Sec. III.2); for the sake of simplicity, we provide an overview of these STA results using the notation that corresponds to 1D systems.
III.1 Invariant-based inverse engineering and the Lewis-Leach family of Hamiltonians
Inverse engineering based on LRIs - the most widely used subclass of STA - have by now become one of the preferred approach for modeling coherent atom transport. The underlying transport theory, advanced in Tor , employs a particular family of quadratic-in-momentum invariants Lew (a). According to that theory, the idealized harmonic-trapping case and that of an arbitrary (anharmonic) trapping potential entail somewhat different treatments. To be more specific, the perfect transport in the latter case requires one, in principle, to utilize compensating forces in the comoving reference frame.
A dynamical invariant of a time-dependent Hamiltonian $H(t)$ is any operator $I(t)$ that satisfies the equation
$\frac{\partial}{\partial t}I(t)+\left[H(t),I(t)\right]=0\>.$ (11)
[…]
is the Lewis-Riesenfeld phase STA . Consequently, the general solution of the TDSE for $H(t)$ can in this case succinctly be written as a weighted sum of transport modes, i.e.
$|\Psi(t)\rangle=\sum_{n}C_{n}\>e^{i\theta^{\textrm{LR}}_{n}(t)}|\Phi_{n}(t)\rangle\>,$ (13)
where $C_{n}$ are constants.
The most important family of Hamiltonians employed in modeling atom transport are those of the Lewis-Leach type Lew (b). These Hamiltonians are given by
$H(t)=\frac{p^{2}}{2m}+V(x,t)\>,$ (14)
where - in the most general case - the potential $V(x,t)$ reads
$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\alpha(t)}{\rho(t)}\right)\>,$ (15)
where $\rho(t)$, $\alpha(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor . In particular, the functions $\rho(t)$ and $\omega(t)$ ought to satisfy the nonlinear second-order ordinary differential equation
$\ddot{\rho}+\omega^{2}(t)\rho=\frac{\omega^{2}_{0}}{\rho^{3}}\>,$ (16)
which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\alpha(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
$\ddot{\alpha}+\omega^{2}(t)\alpha=\frac{F(t)}{m}\>.$ (17)
The quadratic-in-momentum dynamical invariant corresponding to the Hamiltonian in Eq. (14), up to an immaterial additive constant, has the form Tor
$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\alpha})-m\dot{\rho}(x-\alpha)]^{2}$
$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\alpha}{\rho}\right)^{2}+U\left(\frac{x-\alpha}{\rho}\right)\>.$
As it turns out, for a broad class of transport problems one can take $\rho(t)\equiv 1$ and $\omega(t)\equiv\omega_{0}$, effectively rendering Eq. (16) irrelevant; thus, in those cases Eq. (17) remains the only auxiliary equation.
III.2 Harmonic transport: STA treatment
Having reviewed the fundamentals of invariant-based inverse engineering and its application in the special case of the Lewis-Leach-type Hamiltonians [cf. Eq. (15)], it is now pertinent to specialize our discussion to the important case of harmonic transport. In the following, we recapitulate all the relevant general results required to obtain STA transport trajectories using LRI-based inverse engineering.
In transport problems that involve a rigid harmonic oscillator driven by the “transport function” $q_{0}(t)$ – corresponding to $F(t)=m\omega^{2}_{0}q_{0}(t)$ in Eq. (15) – the Eq. (17) adopts the form characteristic of a forced harmonic oscillator and $\alpha(t)$ can be thought of as a classical-particle trajectory $q_{c}(t)$ that satisfies $\ddot{q}_{c}+\omega_{0}^{2}[q_{c}-q_{0}(t)]=0$. The relevant dynamical invariant in the case of harmonic transport – a special case of the general quadratic-in-momentum invariant in Eq. (III.1) – is given by
$I(t)=\frac{1}{2m}\big[p-m\dot{q}_{c}(t)]^{2}\\ +\frac{m}{2}\>\omega^{2}_{0}\left[x-q_{c}(t)\right]^{2}\>.$ (19)
[…]
$\Psi_{n}(x,t)=e^{i\theta^{\textrm{LR}}_{n}(t)}\Phi_{n}(x,t)\equiv e^{i\theta^{\textrm{LR}}_{n}(t)}e^{\frac{i}{\hbar}\>m\dot{q}_{c}(t)x}\phi_{n}(x-q_{c})\>,$ (20)
[…]
$\theta^{\textrm{LR}}_{n}(t)=-i\left(n+\frac{1}{2}\right)\omega_{0}t+\frac{m}{2}\>\int_{0}^{t}\dot{q}_{c}^{2}(t^{\prime})\>dt^{\prime}\>.$ (21)
In order to ensure that $I(t)$ and $H(t)$ are compatible at initial and final times, one sets $[I(t),H(t)]=0$ for $t=0$ and $t=t_{f}$. This translates into the following boundary conditions for the auxiliary function $q_{c}(t)$:
$q_{c}(0)$ $=$ $0\>,\qquad\qquad\>\>q_{c}(t_{f})=d\>,$
$\dot{q}_{c}(0)$ $=$ $\ddot{q}_{c}(0)=0\>,\quad\dot{q}_{c}(t_{f})=\ddot{q}_{c}(t_{f})=0\>.$ (22)
The principal idea in modeling harmonic transport is to first chose a functional form of $q_{c}(t)$ that satisfies these last boundary conditions – typically a polynomial – and then obtain $q_{0}(t)$ through inverse engineering, i.e. from the above forced-harmonic-oscillator equation.
It is important to stress that, in addition to those in Eq. (III.2), one can enforce further boundary conditions on $q_{c}(t)$ – for example, by setting higher-order time derivatives of $q_{c}(t)$ to zero for $t=0,\>t_{f}$. In this manner, one can construct several different auxiliary functions $q_{c}(t)$ for the same transport problem (e.g. polynomials of varying degrees) and compute the corresponding STA transport trajectories $q_{0}(t)$. This non-uniqueness of the resulting trajectory can be exploited to choose a trajectory that satisfies to the largest extent possible certain additional requirements.
IV Dynamical-lattice trajectories
In what follows, we apply the general STA results reviewed in Sec. III to coherent atom transport in the adjustable family of 2D optical lattices under consideration. We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice. We do so within the framework of a single-atom Hamiltonian with the harmonically approximated potential of Eq. (7).
In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$. Here $\bm{\lambda}=\left(\bm{\lambda}_{\mathrm{x}},\bm{\lambda}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\lambda}_{\mathrm{x}}\equiv[\lambda^{(1)}_{\mathrm{x}},\ldots,\lambda^{(6)}_{\mathrm{x}}$] and $\bm{\lambda}_{\mathrm{y}}\equiv[\lambda^{(1)}_{\mathrm{y}},\ldots,\lambda^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively. The latter are assumed to satisfy the following conditions:
$q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
$q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{y}}\>$. (23)
The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\lambda}_{0}\equiv\left(\bm{\lambda}_{\mathrm{x},0},\bm{\lambda}_{\mathrm{y},0}\right)^{\intercal}$, with components $\lambda^{(j)}_{\mathrm{x},0}$ and $\lambda^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$. Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y},0};\>t)$.
We now proceed to apply the LRI-based transport theory to the approximate system Hamiltonian
$H_{\textrm{L},0}(t)=-\frac{\hbar^{2}}{2m}\>\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+V_{\mathrm{L}}(\mathbf{r}-\mathbf{q}_{0}(t))\>,$ (24)
i.e. a single-atom Hamiltonian in which $V(\mathbf{r})\equiv V(x,y)$ is given by the simplified lattice potential $V_{\mathrm{L}}$ of Eq. (7). This approximated potential reads
$V_{\mathrm{L}}\left[\mathbf{r}-\mathbf{q}_{0}(t)\right]=-V_{\mathrm{d,0}}+ma_{\mathrm{x}}\big[x-q_{0}^{\mathrm{x}}(t)\big]$
$+$ $\frac{m}{2}\big\{\omega_{\mathrm{x}}^{2}\big[x-q_{0}^{\mathrm{x}}(t)\big]^{2}+\omega_{\mathrm{y}}^{2}\big[y-q_{0}^{\mathrm{y}}(t)\big]^{2}\big\}\>.$
It is important to stress that – owing to the separability of this last potential – displacements in the $x$- and $y$ directions can be treated independently. Furthermore, upon restricting the above potential to one direction, i.e. making the replacement $V_{\mathrm{L}}\rightarrow V^{\mathrm{x}}_{\mathrm{L}}\left(x-q_{0}^{\mathrm{x}}(t)\right)\>\delta_{x,u}+V^{\mathrm{y}}_{\mathrm{L}}\left(y-q_{0}^{\mathrm{y}}(t)\right)\>\delta_{y,u}+V_{\mathrm{d,0}}$ (where u=x,yu=x,\,y), this potential (up to an additive constant) adopts the characteristic Lewis-Leach form [cf. Eqs. (14) and (15)].
In the problem under consideration, we have a Lewis-Leach-type Hamiltonian with $U=0$ and $\rho(t)\equiv 1$. As it relates to the remaining time-dependent parameters of the general potential in Eq. (15), the following identifications can be made. In the case of atom transport in the $x$ direction, i.e. the potential $V_{\mathrm{L}}^{\mathrm{x}}$, we set $\omega_{\mathrm{x}}(t)=\omega_{\mathrm{x}}$ and
$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0,x}(t)-a_{\mathrm{x}}\right]\>.$ (26)
Analogously, for atom transport along the $y$ direction, i.e. potential $V_{\mathrm{L}}^{\mathrm{y}}$, we have $\omega_{\mathrm{y}}(t)=\omega_{\mathrm{y}}$ and
$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0,y}(t)\>.$ (27)
In keeping with the standard practice Tor , we also add to $V_{\mathrm{L}}^{\mathrm{x}}$ and $V_{\mathrm{L}}^{\mathrm{y}}$ the irrelevant time-dependent global terms $m\omega^{2}_{\mathrm{x}}q^{2}_{0,x}(t)$ and $m\omega^{2}_{\mathrm{y}}q^{2}_{0,y}(t)$, respectively, which do not yield any force. Accordingly, the auxiliary equations of the forced-harmonic-oscillator type here adopt the form
$\ddot{q}_{\mathrm{c,x}}(t)+\omega_{\mathrm{x}}^{2}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]=-a_{\mathrm{x}}\>,$ (28)
$\ddot{q}_{\mathrm{c,y}}(t)+\omega_{\mathrm{y}}^{2}\left[q_{\mathrm{c}}^{\mathrm{y}}(t)-q_{0}^{\mathrm{y}}(t)\right]=0\>,$ (29)
where $\mathbf{q}_{\mathrm{c}}(t)\equiv\left[q_{\mathrm{c}}^{\mathrm{x}}(t),q_{\mathrm{c}}^{\mathrm{y}}(t),0\right]^{\intercal}$ is the corresponding 2D classical-particle trajectory.
One peculiar aspect of these last auxiliary equations is the presence of $a_{x}$ on the right-hand-side of Eq. (28), resulting from the terms linear in $x$ in the approximated potentials $V_{\mathrm{L}}$ [cf. Eq. (7)]. The presence of $a_{x}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor . It should be noted that Eq. (28) can be reduced to the form of Eq. (29) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{x}/\omega_{\mathrm{x}}^{2}$. In this manner, the problem of atom transport in the $x$ direction, equally like the one in the $y$-direction, is reduced to a well-known type of atom-transport problems solvable via inverse engineering Tor .
Figure 4: Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
The boundary conditions for $q_{0}^{\mathrm{x}}(t)$ and $q_{0}^{\mathrm{y}}(t)$ – along with those for for $q_{\mathrm{c}}^{\mathrm{x}}(t)$ and $q_{\mathrm{c}}^{\mathrm{y}}(t)$ consistent with Eqs. (28) and (29) – motivate one to seek the general solution for the path of the potential minimum in the form of the ninth-degree polynomial Hau
$q_{0,u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (30)
where $d_{u}$ is the transport distance. From the relevant initial/boundary conditions Hau , the following results for constants $b_{n,u}$ are obtained:
$b_{3,u}$ $=$ $2520\>(t_{\mathrm{f}}\omega_{u})^{-2}\quad,\quad b_{4,u}=-12600\>(t_{\mathrm{f}}\omega_{u})^{-2}\>,$
$b_{5,u}$ $=$ $22680\>(t_{\mathrm{f}}\omega_{u})^{-2}+126\>,$
$b_{6,u}$ $=$ $-17640\>(t_{\mathrm{f}}\omega_{u})^{-2}-420\>,$ (31)
$b_{7,u}$ $=$ $5040\>(t_{\mathrm{f}}\omega_{u})^{-2}+540\>,$
$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\>.$
The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0,x}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig. 4.
V Calculation of atomic dynamics
In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA-based dynamical-lattice trajectories [cf. Sec. IV]. This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix A).
The crucial prerequisite for computing single-atom dynamics in the problem at hand amounts to evaluating the ground-state wave-function of the relevant optical-lattice potential. Within the general framework of the atom-transport problem (cf. Sec. II.1), that wave function represents the initial ($t=0$) atomic wave packet, i.e. the initial condition for single-atom transport. This wave function can be obtained to requisite accuracy using the plane-wave-expansion method (for details, see, e.g., Sto (b)), which entails Fourier expansions of the underlying lattice potential and the periodic part of single-particle Bloch wave functions in terms of reciprocal-lattice vectors. Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix B). Given that in this study we employ the FSOM for the dynamics part, this last scheme appears as the more convenient choice.
Given the character of the problem that we are concerned with, it is convenient to switch from the lab frame to the comoving frame via a generalized Galilean transformation [cf. Eq. (2)]. In this manner, the lab-frame single-atom TDSE
$i\hbar\>\frac{\partial}{\partial t}\Psi(\mathbf{r},t)=\left[\frac{\hat{\mathbf{p}}^{2}}{2m}+U_{\mathrm{L}}(\mathbf{r}-\mathbf{q}_{0}(t))\right]\Psi(\mathbf{r},t)\>,$ (32)
is transformed into its counterpart in the comoving frame
$i\hbar\>\frac{\partial}{\partial t}\Phi(\mathbf{r},t)$ $=$ $\Big[\frac{\hat{\mathbf{p}}^{2}}{2m}+\frac{m}{2}\>\dot{\mathbf{q}}_{0}^{2}(t)+m\ddot{\mathbf{q}}_{0}\cdot\mathbf{q}_{0}$ (33)
$+$ $m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}(t)+U_{\mathrm{L}}(\mathbf{r})\Big]\Phi(\mathbf{r},t)\>.$
It is important to underscore that the terms $m\dot{\mathbf{q}}_{0}^{2}(t)/2$ and $m\ddot{\mathbf{q}}_{0}(t)\cdot\mathbf{q}_{0}(t)$ in Eq. (33) merely lead to time-dependent global phase factors; the latter are immaterial for the atom-transport problem under consideration and can, accordingly, hereafter be safely neglected. Consequently, the total potential that a transported atom experiences in the comoving frame is given by
$W_{\mathrm{L}}(\mathbf{r},t)\equiv U_{\mathrm{L}}(\mathbf{r})+m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}(t)\>.$ (34)
[…]
$\Psi(\bm{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\bm{r})\frac{\delta t}{2}\right]T_{\mathbf{r},\mathbf{p}}(\delta t)$ (35)
$\times$ $\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)+\mathcal{O}(\delta t^{3})\>,$
where $\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ stands for the operator
$\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)\equiv\exp\left[-\frac{i}{\hbar}\left(\frac{\hat{\mathbf{p}}^{2}}{2m}\>\delta t+m\hat{\mathbf{r}}\cdot\int_{t}^{t+\delta t}\ddot{\mathbf{q}}_{0}(t^{\prime})\mathrm{d}t^{\prime}\right)\right]\>.$ (36)
By making use of the Baker-Campbell-Hausdorff formula and its special case known as the Weyl identity Gil , $\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ can be recast as (for a detailed derivation, see Appendix C in HauSto )
$\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ $=$ $\exp\left[-\frac{i}{\hbar}\frac{\hat{\mathbf{p}}^{2}}{2m}\>\delta t\right]\,\exp\left[-\frac{i}{2\hbar}\>\hat{\mathbf{p}}\cdot\delta\dot{\mathbf{q}}_{0}(t)\>\delta t\right]\,$ (37)
$\times$ $\exp\left[-\frac{i}{\hbar}\>m\hat{\mathbf{r}}\cdot\delta\dot{\mathbf{q}}_{0}(t)\right]\>,$
where $\delta\dot{\mathbf{q}}_{0}(t)\equiv\dot{\mathbf{q}}_{0}(t+\delta t)-\dot{\mathbf{q}}_{0}(t)$. It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the RHS of the last equation, analogously to the dropped terms in Eq. (33) above.
With the aid of an analog of Eq. (A) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (37), switching also to the coordinate representation of the momentum operator [i.e. $\hat{\mathbf{p}}\rightarrow(\hbar/i)\nabla$], we arrive at the FSOM-based (second-order) time-stepping scheme for the atom-transport problem at hand:
$\Phi(\mathbf{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\mathbf{r})\frac{\delta t}{2}\right]\mathrm{F}^{-1}\Bigg[e^{-i\frac{\hbar k^{2}}{2m}\delta t}$ (38)
$\times$ $\exp\left[-\frac{i}{2}\>\mathbf{k}\cdot\delta\dot{\mathbf{q}}_{0}\delta t\right]\mathrm{F}\Bigg\{\exp\left[-\frac{i}{\hbar}\>m\mathbf{r}\cdot\delta\dot{\mathbf{q}}_{0}\right]$
$\times$ $\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\mathbf{r})\frac{\delta t}{2}\right]\Phi(\mathbf{r},t)\Bigg\}\Bigg]+\mathcal{O}(\delta t^{3})\>.$
In order to alleviate the computational burden in the problem at hand, we exploit the discrete translational symmetry of the system and restrict ourselves to the displacement of one unit cell of the original optical lattice. The appropriate choice of this unit cell has to be consistent with the transport direction. For instance, in order to evaluate transport in the $x$- and $y$ directions we make use of a rectangular unit cell (indicated by white dashed lines in Fig. 2).
Yet another problem-specific feature that allows one to further reduce the level of computational difficulty in the problem at hand is closely related to its description in the comoving frame. Namely, the form of the relevant TDSE [cf. Eq. (33)] immediately implies that there is no need to compute the potential term after each time step, but only the correction term $m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}$ [cf. Eq. (36)] that depends on the acceleration $\ddot{\mathbf{q}}_{0}$ of the potential minimum.
The spatial Fourier transformation is carried out using the FFT algorithm (cf. Appendix ANRc . To be more specific, we make use of a discrete three-dimensional $N_{x}\times N_{y}\times N_{z}$ grid with $N_{x}=200$, $N_{y}=300$, and $N_{z}=100$ points in the computational window whose dimensions (in units of the respective characteristic lenghtscales $l_{\mathrm{x}}$, $l_{\mathrm{y}}$, and $l_{\mathrm{z}}$) are $100$, 100100, and $500$, respectively. For the propagation of the TDSE in the comoving frame [cf. Eq. (33)], we make use of the adaptive approach NRc with a maximal relative error of $10^{-4}$; the latter could – strictly speaking – require additional (smaller in size) substeps in time. Yet, it turned out to be sufficient to take $N_{t}$ between $20$ and $100$ in all our runs.
VI Results and discussion
In what follows, we quantify and discuss the efficiency of STA-based coherent single-atom transport (cf. Sec. IV above) in various optical lattices belonging to the lattice family under consideration. The main figure of merit used for this purpose is the transport fidelity, defined as
$\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}\>.$ (39)
Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix B) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec. V above and Appendix A).
Figure 5: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ and potential depths using the STA scheme. The ratio of the lattice depths corresponds to that of a honeycomb lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,600,200)E_{R}$, $(400,1200,400)E_{R}$, and $(800,2400,800)E_{R}$, respectively.
Figure 6: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a square lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
Figure 7: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a linear lattice (1D chains). Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
Figure 8: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a dimerized lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,400,200)E_{R}$, $(400,800,400)E_{R}$, and $(800,1600,800)E_{R}$, respectively.
The obtained results for the transport fidelity in the optical lattices of the adjustable family under consideration are depicted in Figs. 5-8. One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{f}$. This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$ that is proportional to the lattice depth Hau the atom being transported cannot be confined by the optical-lattice potential. This happens as a consequence of the disappearance of the minima of the “tilted” lattice potential experienced by the atom in its comoving (non-inertial) reference frame; compared to the potential in the lab frame, the comoving-frame potential has an additional term linear in the transport coordinate Hau . This characteristic behavior for short transport times is a generic property of coherent atom transport in optical lattices, i.e. it is not special for the concrete family of lattices being considered here.
The aforementioned breakdown of transport occurs when the maximal acceleration of an atom during transport - hereafter denoted by $|\tilde{a}_{\textrm{max}}|$ - becomes larger than $|a_{\textrm{max}}|$. It should be borne in mind that the actual value of $|\tilde{a}_{\textrm{max}}|$ depends on the chosen dynamical-lattice trajectory (i.e. the path of the potential minimum); yet, quite generally it holds that $|\tilde{a}_{\textrm{max}}|\propto t_{f}^{-2}$. It immediately follows from this last result that for deeper lattices – i.e., for a larger value of $|a_{\textrm{max}}|$, which is proportional to the lattice depth – the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached becomes larger than $|a_{\textrm{max}}|$ for sufficiently short total time $t_{f}$. In other words, for deeper lattices the breakdown of transport (i.e., the abrupt decrease of the transport fidelity) takes place for shorter times $t_{f}$.
Our obtained numerical results for the transport fidelity [cf. Eq. (39)] are displayed in Figs. 5-8 for honeycomb, square, linear (1D chain), and dimerized lattices, respectively, and two different transport distances ($d=100\>l_{x}$ and $d=400\>l_{x}$). These results are consistent with the above general features of coherent atom transport in optical lattices. In other words, they clearly manifest breakdown of transport at short transport times. On the other hand, in the opposite limit of very long transport times $t_{\mathrm{f}}$ the results coincide with the well-known adiabatic limit, i.e. essentially perfect atom transport ($\mathcal{F}\approx 1$) for $t_{\mathrm{f}}\gg\omega_{\mathrm{j}}^{-1}$ ($j=x,y$) is achieved.
In particular, in the case of the honeycomb lattice, with the parameter values chosen as in Fig. 5, the respective transport times $t_{f}$ corresponding to the onset of the breakdown of atom transport are in the range between $7\>T_{\mathrm{x}}$ and $8\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, as well as in the range between $16\>T_{\mathrm{x}}$ and $17.5\>T_{\mathrm{x}}$ for the transport distance $d=400\>l_{x}$.
In the case of the square lattice, which corresponds to Fig. 6, the breakdown of atom transport takes place at times in the range between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$. These results are clearly in agreement with the anticipated trend of decreasing with the increase of the lattice depth.
For the linear (1D-chain) lattice that corresponds to Fig. 7, the transport breakdown takes place when $t_{f}$ is between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
Finally, in the case of the dimerized lattice with the parameter values chosen as in Fig. 8, the onset of the transport breakdown occurs for $t_{f}$ between $10\>T_{\mathrm{x}}$ and $13\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $14\>T_{\mathrm{x}}$ and $18\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
VII Summary and Conclusions
To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices. We first specified concrete arrangements of acousto-optic modulators that give rise to the moving-lattice effect in these systems in an arbitrary direction. We subsequently designed the corresponding dynamical-lattice trajectories using shortcuts to adiabaticity (STA). Through numerical propagation of the relevant time-dependent Schrödinger equation – using the previously computed STA-based dynamical-lattice trajectories – we calculated the efficiency ofhe resulting coherent atom transport as described by the atom-transport fidelity.
We demonstrated that the STA approach enables fast atom transport in the considered family of optical lattices, for a wide range of lattice depths and transport directions. In addition, we analyzed the implications of the atom-transport processes for neutral-atom quantum computing based on collisional two-qubit entangling gates.
By contrast to previous theoretical investigations of coherent atom transport, based mostly on simplified scenarios, in the present work we addressed this phenomenon based on the full, anharmonic optical-lattice potential. Thus, the obtained results can be corroborated in future transport experiments in such lattices. Furthermore, our study also complements well the recent works on quantum-state control of Bose-Einstein condensates in optical lattices Amr . Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems, the first one of which has been undertaken quite recently Hwa .
Acknowledgements.
V.M.S. is grateful to K. Viebahn for calling his attention to the dynamical optical-lattice family of Tarr . This research was supported by the Deutsche Forschungsgemeinschaft (DFG) – SFB 1119 – 236615297.
Appendix A FSOM-based numerical propagation of TDSEs
In the following, we briefly review the essential ingredients of the use of the FSOM for solving TDSEs.
Let us consider Cauchy-type initial-value problems $\partial_{t}f(\mathbf{r},t)=\hat{A}(\mathbf{r},t)f(\mathbf{r},t)$, where $\hat{A}(\mathbf{r},t)$ is an operator and the initial condition is given by $f(\mathbf{r},t)=f_{0}(\mathbf{r})$. The FSOM is the method of choice for solving such problems in cases where $\hat{A}$ can be written as a sum of an operator that can straightforwardly be diagonalized in real space and another one that can be diagonalized in Fourier space. One of the most common examples of such problems is the TDSE Fei ; Ban . The crucial idea facilitating the use of FSOM for solving TDSEs is to approximate the time-evolution operator of the system under consideration by a product of operators that are either diagonal in real space or in Fourier space.
Consider the TDSE that describes the motion of a particle of mass $m$ in the potential $U(\bm{r})$. The underlying single-particle Hamiltonian is given by
$H=-\frac{\hbar^{2}\nabla^{2}}{2m}+U(\bm{r})\>.$ (40)
To approximate the time-evolution operator that corresponds to this last Hamiltonian we invoke the identity (known as the symmetrized Trotter decomposition) Gil
$e^{i(A+B)\delta t}=e^{iA\delta t/2}e^{iB\delta t}e^{iA\delta t/2}+\mathcal{O}(\delta t^{3})\>,$ (41)
which holds for arbitrary Hermitian operators $A$ and $B$. In the special case $A=U(\bm{r})$, $B=-\hbar^{2}\nabla^{2}/(2m)$ of this last identity, we obtain an accurate second-order time-stepping scheme  Pec
$\Psi(\bm{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\exp\left(i\frac{\hbar\nabla^{2}}{2m}\>\delta t\right)$ (42)
$\times$ $\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)+\mathcal{O}(\delta t^{3})$
for the propagation of the wave-function $\Psi(\bm{r},t)$ of a system described by the Hamiltonian in Eq. (40). This approach to evaluating the time evolution – in which the potential part of the total Hamiltonian is applied for a half time step before and after the kinetic part – is often referred to as symmetric splitting (or, alternatively, as Strang splitting Spe ) and its principal advantage compared to the conventional approach is that it reduces the error to $\mathcal{O}(\delta t^{3})$.
The form of Eq. (42) permits independent treatment of the different exponential terms. Therefore, it is pertinent to Fourier-transform the kinetic term, recasting at the same time the RHS of Eq. (42) using the identity
$\exp\left(i\frac{\hbar\nabla^{2}}{2m}\>\delta t\right)\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)$
$=$ $F^{-1}\left[e^{-i\frac{\hbar k^{2}}{2m}\delta t}\>F\left\{\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)\right\}\right]\>,$
where $F[\ldots]$ stands for the Fourier transform of the argument and $F^{-1}[\ldots]$ for its inverse.
By successively applying the time-propagation scheme based on Eq. (42) $N_{t}$ times on an initial wave-function $\Psi(\bm{r},t)$ one obtains a numerical solution of a TDSE at time $t^{\prime}=t+N_{t}\delta t$. In actual numerical implementations of the FSOM, the wave-function is represented on a discrete rectangular spatial grid with $N_{s}$ points and the exact (continuous) Fourier transform is approximated by its discrete counterpart. The computational burden involved in propagating the wave function is dominated by the need to compute its spatial Fourier transformation and its inverse. If carried out using the fast Fourier transform (FFT) algorithm NRc , an elementary step in these transformations entails $\mathcal{O}(N_{s}\log_{2}N_{s})$ operations.
Appendix B Ground-state calculation via the ITE approach
Here we provide a short introduction into the ITE approach, which established itself as one of the most robust computational methods for obtaining ground states of quantum systems And ; Fei . Its relationship to the FSOMa - its counterpart utilized for computing dynamics - bears close analogy to the connection between Lanczos-type diagonalizations and Chebyshev-propagator methods Stojanović and Salom (2019). It is prudent to start our review of the ITE approach using the abstract, representation-independent notation for quantum states.
Consider the problem of obtaining the ground state of a Hamiltonian denoted by $H$ in the following. One starts by expanding a quantum state $|\phi\rangle$ in terms of the eigenstates $\{|\psi_{n}\rangle,\>n=0,1,\ldots\}$ of $H$, i.e. $|\phi\rangle=\sum_{n}c_{n}|\psi_{n}\rangle$, assuming that the sought-after ground state $|\psi_{0}\rangle$ has a nonvanishing contribution to this expansion (i.e. $c_{0}\equiv\langle\psi_{0}|\phi\rangle\neq 0$). Assuming that $|\phi\rangle$ is the initial ($t=0$) state of the system, its states $|\phi(t)\rangle\equiv e^{-iHt/\hbar}|\phi\rangle$ at later times $t>0$ are given by the linear combinations
$|\phi(t)\rangle=\sum_{n}c_{n}\>e^{-iE_{n}t/\hbar}|\psi_{n}\rangle$ (44)
of the time-evolved eigenstates $e^{-iE_{n}t/\hbar}|\psi_{n}\rangle$ of $H$, where $E_{n}$ is the eigenvalue of $H$ that corresponds to the eigenstate $|\psi_{n}\rangle$.
By performing a Wick rotation into the complex plane (i.e. by making the substitution $t\rightarrow\tau=it$), Eq. (44) goes over into
$|\phi(\tau)\rangle=\sum_{n}c_{n}\>e^{-E_{n}\tau/\hbar}|\psi_{n}\rangle\>.$ (45)
Therefore, $|\phi(\tau)\rangle$ is an exponentially-decaying superposition of the eigenstates $|\psi_{n}\rangle$ with the decay rates given by the energy eigenvalues $E_{n}$. Given that the decay rate of the ground state is the smallest one, in the large-$\tau$ limit one arrives at
$|\phi(\tau)\rangle\approx c_{0}\>e^{-E_{0}\tau/\hbar}|\psi_{0}\rangle\>.$ (46)
Therefore, given a trial state $|\phi(\tau=0)\rangle\equiv|\phi\rangle$ with a nonvanishing overlap with the sought-after ground state, the ITE approach yields a monotonously decreasing functional of $\tau$ that converges to the ground-state energy $E_{0}$ of the system as $\tau\rightarrow\infty$, i.e.
$E_{0}=\lim_{\tau\rightarrow\infty}\langle\phi(\tau)|\>H\>|\phi(\tau)\rangle\>.$ (47)
In the atom-transport problem under consideration (cf. Sec. II.1) – and many other problems – one makes use of the coordinate representation, in which a quantum state $|\psi\rangle$ is represented by the wave function $\psi(\mathbf{r})\equiv\langle\mathbf{r}|\psi\rangle$ (with $|\mathbf{r}\rangle$ being the eigenstates of the position operator). The counterpart of Eq. (45) in this representation reads
$\phi(\bm{r},\tau)=\sum_{n}c_{n}e^{-E_{n}\tau/\hbar}\psi_{n}(\bm{r})\>,$ (48)
where $\phi(\bm{r},\tau)\equiv\langle\mathbf{r}|\phi(\tau)\rangle$ and $\psi_{n}(\bm{r})\equiv\langle\mathbf{r}|\psi_{n}\rangle$ are the relevant wave functions. The remaining steps in the application of the ITE approach within this representation are carried out in complete analogy with the above representation-independent expressions (46) and (47).

### v2

@@
-Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices
+Enhanced shortcuts to adiabaticity for coherent atom transport in a family of two-dimensional dynamical optical lattices
 Abstract
-Motivated by the compelling need for coherent atom transport in a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L.
-Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, triangular, honeycomb, dimerized, and checkerboard lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms.
+In view of the compelling need for coherent atom transport as a prerequisite for a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L.
+Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms.
 At the outset, we propose the appropriate arrangements of acousto-optic modulators that give rise to a frequency imbalance between counterpropagating laser beams, thus leading to the dynamical-lattice effect in an arbitrary direction in the lattice plane.
-We subsequently obtain the dynamical-lattice trajectories that enable atom transport using shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type.
-We then quantify the resulting atom dynamics using the transport fidelity computed from the numerical solutions of the relevant time-dependent Schrödinger equation.
-We do so for various choices of the system parameters and transport directions, finding favorable results for the achievable transport times and robustness of the resulting transport to various experimental imperfections.
+We subsequently obtain the dynamical-lattice trajectories that enable atom transport in the lattices under consideration using two classes of control schemes: (i) shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type, and (ii) their modification, known as enhanced STA (eSTA), which is well-suited for the treatment of anharmonic trapping potentials.
+We then quantify the resulting atom dynamics using transport fidelities computed from the numerical solutions of the relevant time-dependent Schrödinger equations.
+By doing so for various choices of the system parameters and transport directions, we demonstrate that – except in the special case of the dimerized lattice – the eSTA method consistently outperforms its STA counterpart, both in terms of the achievable transport times and the robustness of the resulting transport against small variations of optical-lattice depths.
 I Introduction
-Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic ; Ton .
+Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic .
 As a result, coherent atom transport Qic enabled by moving the confining optical trap – either an optical lattice, an array of optical tweezers, or their combination – has lately been attracting considerable attention, on the experimental- and theoretical sides alike.
-While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STAf ; ContRev .
+While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STA ; STA ; STAf ; ContRev .
 Coherent single-atom transport in an optically-trapped system entails a final atomic state that approximates - as closely as possible - the initial state in the comoving reference frame (i.e.
@@
 Owing to the fact that adiabatic control-parameter variations typically render some dynamical properties of the system invariant, inverse-engineering schemes based on Lewis-Riesenfeld invariants (LRIs) Lew (a) constitute the most pervasive subclass of STA quantum-control protocols.
-The early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework.
-However, given that such simplifications typically do not apply in realistic systems - where there is either a significant coupling between the longitudinal and transverse degrees of freedom, or the relevant confining potential is anharmonic Zha subsequent theoretical studies have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hau and double-well  HauSto types.
+Early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework.
+However, given that such simplifications typically do not apply in realistic systems - where either a significant coupling between the longitudinal and transverse degrees of freedom is present, or the relevant confining potential is anharmonic Zha - subsequent theoretical studies Hau ; HauSto have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hic and double-well Sto (b); Hof types.
 Atom transport in such systems is enabled by the dynamical-lattice effect, which is induced by appropriately arranged pairs of acousto-optic modulators (AOMs) that give rise to a frequency imbalance between two counterpropagating laser beams.
-These studies of atom transport in higher-dimensional optical lattices Sto (b); Hof were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials.
+These studies of atom transport in higher-dimensional optical lattices were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials.
 The results obtained using this method are comparable to those corresponding to the quantum-speed limit HauSto .
-In this paper, we aim to demonstrate – using the STA approach – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family of lattices includes square, triangular, dimerized, honeycomb, and checkerboard lattices as its special cases.
-Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie ; these gates originate from exchange interactions between qubits Heu ; Heu (2); Sto (e); Hay .
+In this paper, we aim to demonstrate – using both the STA approach and eSTA-based modification – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases.
+Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie .
 We start by proposing specific arrangements of AOMs that enable the dynamical-lattice effect in an arbitrary direction in the relevant lattice plane.
-We then set out to model the ensuing coherent atom transport by employing STA-type protocols, here based on inverse-engineering techniques involving LRIs.
-Having obtained the STA dynamical-lattice trajectories, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM).
-We then quantify the efficiency of STA-based atomic transport through the corresponding transport fidelity for a broad range of values of the relevant system parameters.
-In addition to finding STA solutions to the atom-transport problem at hand, we devote significant attention to analyzing their sensitivity to errors, i.e.
-their robustness to various experimental imperfections, such as deviations in the control amplitudes, the duration of control, the unavoidable presence of noise, etc.
+We then set out to model the ensuing coherent atom transport by employing two classes of control protocols: (i) STA-type protocols, here based on inverse-engineering techniques involving LRIs, and (ii) protocols based on the eSTA method, which unlike STA-based ones systematically take into account the full optical-lattice potential.
+Having obtained the dynamical-lattice trajectories based on each of these two methods, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM).
+We then quantify and compare the efficiency of STA- and eSTA-based atomic transport through the corresponding transport fidelities for a broad range of values of the relevant system parameters.
+In the manner described above, we show that both approaches – STA and eSTA – allow time-efficient atom transport in the family of optical lattices under consideration.
+We demonstrate that while STA provides somewhat faster transport in shallow lattices (i.e.
+for relatively small lattice depths), eSTA becomes far superior to it for deeper lattices.
+Based on our numerical evaluations, we show that this superiority of the eSTA approach compared to its STA counterpart is more pronounced for shorter transport distances.
+In addition to finding both STA- and eSTA solutions to the atom-transport problem at hand, we devote due attention to analyzing their sensitivity to errors, more precisely their robustness against deviations in optical-lattice depths.
+The obtained results bear out the superior robustness of the eSTA protocols compared to their STA counterparts in the transport problem under consideration, which to a large extent originates from the general assumptions underlying the inner workings of the eSTA approach.
 The remainder of this paper is organized as follows.
@@
 II by providing the basic background of the atom-transport problem at hand, including the Hamiltonian describing such transport and the essential properties of the relevant family of optical lattices; we then specify concrete arrangements of AOMs that enable atom transport in the system under consideration.
-Section III is devoted to a short survey of the general STA-related results of relevance for the problem at hand.
+Section III is devoted to a short survey of the general STA- and eSTA-related results of relevance for the problem at hand.
 In the following, Sec.
-IV, we describe how these general results are utilized for obtaining STA dynamical-lattice trajectories and analyze typical examples thereof.
+IV, we describe how these general results are utilized for obtaining STA and eSTA dynamical-lattice trajectories and analyze typical examples thereof.
 Section V is devoted to the description of our approach for the treatment of atomic dynamics.
 In Sec.
-VI, we discuss the obtained results for the atom-transport fidelities.
+VI, we discuss the results for the atom-transport fidelities obtained using either of the two approaches; we then discuss how these results are affected by the presence of small lattice-depth variations, and, finally, also demonstrate the superiority of our adopted transport schemes to some alternative approaches for modeling coherent atom transport in optical lattices.
 To end with, a short summary of the paper - along with the main conclusions drawn - is provided in Sec.
 VII.
-In Appendix A we briefly review the application of the FSOM for solving TDSEs, while Appendix B contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems.
+In order not to interrupt the flow of the discussion in the main part of this paper, certain cumbersome mathematical derivations are relegated to Appendices A and B.
+For similar reasons, in Appendix C we briefly review the application of the FSOM for solving TDSEs, while Appendix D contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems.
+Finally, in Appendix E we derive a dimensionless set of units that enables more accurate numerical simulations than the original units.
 II System and atom transport
@@
 (1), in Sec.
-IV below we discuss how – using an STA-based method – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration.
+IV below we discuss how – using STA- and eSTA methods – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration.
 In particular, both of our approaches (see Sec.
@@
 II.2 Adjustable family of 2D optical lattices
-In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remainder of this work.
+In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remaining part of this work.
 In order to provide a more concise and clearer explanation of the envisioned setup, we hereafter restrict our discussion to the strong confining potentials in the xyx-y plane; we neglect the minor modulations arising from the ever-present confinement of the system in the $z$ direction, which in the anticipated physical realization is close to being harmonic.
@@
 1), with $I_{\overline{X}}$, $I_{X}$, and $I_{Y}$ being the respective beam intensities.
-In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis (cf.
-Fig.
-1).
+In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis.
 In particular, the interference of two perpendicular beams – namely, beams $X$ and $Y$ – creates a checkerboard lattice with period $\lambda_{\mathrm{L}}/\sqrt{2}$.
@@
 The potential that corresponds to the envisioned adjustable family of 2D optical lattices can be derived starting from the general expression for the light intensity $I(x,y)$ in terms of the electric field $\mathbf{E}(x,y)$ Sho
-$I(x,y)=c\epsilon_{0}\>|\mathbf{E}(x,y)|^{2}\>.$ (3)
+$I(x,y)=c\varepsilon_{0}\>|\mathbf{E}(x,y)|^{2}\>.$ (3)
 By making use of this last expression, one finds that the intensity pattern corresponding to the checkerboard lattice is given by
 $I_{1}(x,y)=$ - $I_{X}\,\sin^{2}(k_{\mathrm{L}}x)-I_{Y}\,\cos^{2}(k_{\mathrm{L}}y)$
-- $2\alpha\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
-where $\phi$ corresponds to the relative phase between the two arms, while $\alpha$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting.
+- $2\gamma\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
+where $\phi$ corresponds to the relative phase between the two arms, while $\gamma$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting.
 The possible sources of experimental imperfections are, for instance, imperfect coherence and mode-mismatching.
-In what follows, we will restrict our analysis to high-contrast regimes of $\alpha=0.9$, which is readily achievable in experimental realizations Tarr .
+In what follows, we will restrict our analysis to high-contrast regimes of $\gamma=0.9$, which is readily achievable in experimental realizations Tarr .
 Figure 1: Schematic of a setup that enables the adjustable optical-lattice formed from three folded retro-reflected beams.
@@
 - $U_{\overline{X}}\,\cos^{2}(k_{\mathrm{L}}x+\theta/2)$
-- $2\alpha\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
+- $2\gamma\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
 where $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ are the three single-beam lattice depths (proportional to the laser-beam intensities).
@@
 2.
-Among them, the following four lattice structures can be realized: square, dimerized, honeycomb, and 1D chains (cf.
-Fig.
-2).
+Among them, the following four lattice structures can be realized: honeycomb, triangular, dimerized, and square.
+These four lattice types will be utilized in the following to test the performance of the eSTA scheme compared to its STA counterpart in modeling coherent atom transport.
 Figure 2: Density plots of the optical potential for the (a) square, (b) dimerized, (c) honeycomb, and (d) 1D-chains, optical lattices, obtainable by adjusting the different lattice depths present in the expression for $U_{\mathrm{L}}(x,y)$ [cf.
@@
 For each lattice structure, the corresponding unit cell is indicated by the dashed white line.
-The point $(x_{0},y_{0})=(0,0)$, around which the harmonic approximation of the full potential is developed, lies in the center od the unit cell.
+The point $(x_{0},0)$, around which the harmonic approximation of the full potential is developed, lies in the center of the unit cell.
 Having analyzed the full, anharmonic optical-lattice potential of the system at hand [cf.
@@
 III.2 and IV below).
-Given that the adjustable optical-lattice potential in Eq.
-(II.2) has minima around $(x_{0},y_{0})=(0,0)$ for all previously discussed configurations (cf.
-Fig.
-2), we will approximate the full potential around the origin.
-We do so assuming that $k_{\mathrm{L}}|x|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$.
-Based on these assumptions, we straightforwardly obtain
+Given that the adjustable optical-lattice potential in Eq. (II.2) has a minimum at the point $(x_{0},y_{0})$ with
+$x_{0}$ $=\frac{1}{k_{\mathrm{L}}}\arcsin\left(\gamma\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}-U_{X}}\right)\>,$ (7)
+$y_{0}$ $=0\>,$ (8)
+for all previously discussed lattice configurations (see Fig.
+2) we will approximate the full potential around their respective minima.
+We do so assuming that $k_{\mathrm{L}}|x-x_{0}|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$.
+Based on these last assumptions, it is straightforward to obtain
 $V_{\mathrm{L}}=-V_{\mathrm{d},0}+ma_{\mathrm{x}}x+\frac{m}{2}\left(\omega_{\mathrm{x}}^{2}x^{2}+\omega_{\mathrm{y}}^{2}y^{2}\right)\>,$ (9)
 where $V_{\mathrm{d},0}$ depends on $\theta$ and is given by
-$V_{\mathrm{d},0}=U_{Y}+U_{\overline{X}}\cos^{2}\left(\frac{\theta}{2}\right)\>.$ (8)
-The squared frequencies in Eq. (7) are given by
-$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\>,$ (9)
-$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\>,$
-while the expression for the factor $a_{\mathrm{x}}$, which has dimensions of acceleration, reads
-$a_{\mathrm{x}}=\frac{U_{\overline{X}}k_{\mathrm{L}}}{m}\left(\sin\theta-2\alpha\>\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}}\cos\phi\right)\>.$ (10)
+$V_{\mathrm{d},0}=$ $U_{X}\sin^{2}(kx_{0})+\frac{U_{\overline{X}}}{2}\left[1+\cos(2kx_{0})\cos\theta\right]$ (10)
+$+U_{Y}+2\gamma\sqrt{U_{X}U_{Y}}\sin(kx_{0})\cos\phi\>.$
+The explicit expressions for the squared frequencies in Eq. (9) read
+$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\cos^{2}(kx_{0})\>,$ (11)
+$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\left[1+\sqrt{\frac{U_{X}}{U_{Y}}}\>\gamma\sin(k_{\mathrm{L}}x_{0})\cos\phi\right]\>,$
+while the factor $a_{\mathrm{x}}$, with dimensions of acceleration, is given by
+$a_{\mathrm{x}}=-\omega_{\mathrm{x}}^{2}x_{0}\>.$ (12)
 Therefore, the approximated potential has a term linear in $x$.
@@
 (9), of interest for our further discussion, is that – by contrast to the total lattice potential, which is nonseparable in the xyx-y plane (i.e.
-it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by this approximated potential.
+it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by the approximated potential.
 II.3 AOM configurations for atom transport
-To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$, $y$, and diagonal ($r$) directions.
-To be able to appreciate the specific configurations of AOMs that give rise to the dynamical-lattice effect in the system at hand (see Fig.
-3 below), it is helpful to first recall how the use of an AOM (also known as a Bragg cell) gives rise to the dynamical-lattice effect.
-AOMs lead to a shift of an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal.
-Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to Bragg diffraction.
-Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua .
-In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs.
-While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wave-numbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively.
-This imbalance engenders a dynamical (moving) optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wave number HauSto .
+To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$ and $y$ directions (for an illustration, see Figs.
+1 and 3).
 Figure 3: Adapted schematic of a setup that enables atom transport for the adjustable optical-lattice.
@@
 The red dot indicates the center of the optical-lattice potential.
+To be able to appreciate the proposed configurations of AOMs, it is helpful to first recall how the use of an AOM gives rise to the dynamical-lattice effect.
+AOMs lead to a shift an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal.
+Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to diffraction.
+Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua .
+In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs.
+While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wavenumbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively.
+This imbalance engenders a moving optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wavenumber HauSto .
 Having described the general procedure for incorporating AOMs into the system to enable atom transport, we now discuss how this approach can be implemented in the setup for the adjustable potential shown in Fig.
@@
 Currently, a single AOM is employed to tune the phase shift $\theta$ present in the optical-lattice potential of Eq.
-II.2.
+(II.2).
 However, it is important to clarify that this particular AOM does not participate in the shifting procedure.
@@
 For ideal diagonal transport, the frequency differences in both arms must be tuned identically.
-III STA: survey of results relevant for atom transport
+III STA and eSTA: survey of results relevant for atom transport
 To set the stage for modeling coherent atom transport in the family of optical lattices under consideration using the STA- and eSTA methods (see Sec.
@@
 III.2); for the sake of simplicity, we provide an overview of these STA results using the notation that corresponds to 1D systems.
+Finally, we discuss the basic aspects of the eSTA method, explaining in detail how an eSTA control scheme can be obtained from its STA counterpart that is exact for a simplified Hamiltonian of the system (Sec.
+III.3).
 III.1 Invariant-based inverse engineering and the Lewis-Leach family of Hamiltonians
@@
 where - in the most general case - the potential $V(x,t)$ reads
-$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\alpha(t)}{\rho(t)}\right)\>,$ (15)
-where $\rho(t)$, $\alpha(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor .
+$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\beta(t)}{\rho(t)}\right)\>,$ (17)
+where $\rho(t)$, $\beta(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor .
 In particular, the functions $\rho(t)$ and $\omega(t)$ ought to satisfy the nonlinear second-order ordinary differential equation
 $\ddot{\rho}+\omega^{2}(t)\rho=\frac{\omega^{2}_{0}}{\rho^{3}}\>,$ (18)
-which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\alpha(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
-$\ddot{\alpha}+\omega^{2}(t)\alpha=\frac{F(t)}{m}\>.$ (17)
+which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\beta(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
+$\ddot{\beta}+\omega^{2}(t)\beta=\frac{F(t)}{m}\>.$ (19)
 The quadratic-in-momentum dynamical invariant corresponding to the Hamiltonian in Eq. (16), up to an immaterial additive constant, has the form Tor
-$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\alpha})-m\dot{\rho}(x-\alpha)]^{2}$
-$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\alpha}{\rho}\right)^{2}+U\left(\frac{x-\alpha}{\rho}\right)\>.$
+$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\beta})-m\dot{\rho}(x-\beta)]^{2}$
+$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\beta}{\rho}\right)^{2}+U\left(\frac{x-\beta}{\rho}\right)\>.$
 As it turns out, for a broad class of transport problems one can take $\rho(t)\equiv 1$ and $\omega(t)\equiv\omega_{0}$, effectively rendering Eq.
@@
 This non-uniqueness of the resulting trajectory can be exploited to choose a trajectory that satisfies to the largest extent possible certain additional requirements.
+III.3 Basics of the eSTA method
+While STA-based control protocols have established themselves in a variety of quantum systems STA , their modification – which became known as eSTA – has been proposed Whi (a).
+The motivation behind this enhanced version of the STA approach, which is inspired in part by optimal-control methods Whi (a), is to facilitate the development of efficient control schemes for systems not directly amenable to an STA-type treatment.
+The principal idea of eSTA is to approximate the original Hamiltonian of such a system by a simpler one that admits a straightforward design of an STA-based protocol.
+Assuming that this last protocol is nearly optimal even when applied to the original Hamiltonian, its eSTA counterpart is constructed using a gradient expansion in the control-parameter space.
+This crucial assumption underscores the semi-heuristic character of the eSTA method, which – in principle – is not guaranteed to outperform STA Whi (a).
+Yet, it has already been demonstrated that eSTA is superior to STA in several classes of quantum-control problems HauSto ; Odell (1).
+The initial step that the application of the eSTA method to a system described by the Hamiltonian $H_{\textrm{S}}$ entails amounts to obtaining an STA solution for a “close,” approximated Hamiltonian $H_{0}$; the latter corresponds to the value $\bm{\zeta}_{0}$ of the vector parameter $\bm{\zeta}\in\mathbb{R}^{n}$ Whi (a).
+It is assumed that a value $\mu_{\textrm{S}}$ of a scalar parameter $\mu$ can be found such that a series expansion
+$H_{\textrm{S}}=\sum_{k=0}^{\infty}\mu_{\textrm{S}}^{k}\,H^{(k)}\>,$ (25)
+has $H_{0}$ as its zeroth-order term [i.e. $H^{(0)}\equiv H_{0}$].
+In order to find an eSTA solution for $H_{\textrm{S}}$ based on a previously found STA solution for $H_{0}$, the control vector $\bm{\zeta}_{\mathrm{S}}$ corresponding to $H_{\textrm{S}}$ can be written as a sum of the STA control vector $\bm{\zeta}_{0}$ and an auxiliary control vector $\bm{\alpha}$, i.e.
+$\bm{\zeta}_{\mathrm{S}}=\bm{\zeta}_{0}+\bm{\alpha}$.
+The value of $\bm{\alpha}$ that corresponds to the desired eSTA solution, will be denoted by $\bm{\epsilon}$ in the following.
+As stated above, the principal assumption underlying the eSTA scheme is that the STA-based protocol corresponding to the Hamiltonian $H_{0}$ is near-optimal even when applied to the full Hamiltonian $H_{\textrm{S}}$ Whi (a).
+Another important assumption within the eSTA approach is that the deviation of the fidelity $\mathcal{F}$ from its maximal value is quadratic in the difference $\alpha-\epsilon$, i.e.
+the fidelity can be written in the form Whi (a)
+$\mathcal{F}\left(\mu_{\textrm{S}},\bm{\zeta}_{0}+\alpha\frac{\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})}{\rVert\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\rVert}\right)\approx 1-c\left(\alpha-\epsilon\right)^{2}\>,$ (26)
+where $\epsilon\equiv\rVert\bm{\epsilon}\rVert$, $\alpha\equiv\rVert\bm{\alpha}\rVert$, and $c$ is a positive constant.
+The above assumptions and a Taylor expansion of the left-hand-side of Eq. (26) around $\epsilon=\alpha$, straightforwardly lead to  Whi (a)
+$\bm{\epsilon}\approx\frac{2\left[1-\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{\textrm{S}})\right]\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})}{\rVert\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\rVert^{2}}\>.$ (27)
+To second order in $\mu_{\textrm{S}}$ the fidelity is given by Whi (a)
+$\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{\textrm{S}})\approx 1-\frac{1}{\hbar^{2}}\sum_{n=1}^{\infty}|G_{n}|^{2}\>,$ (28)
+where $G_{n}$ is an auxiliary scalar function, given in terms of the transport modes of $H_{0}$ [cf. Sec. III.1] by
+$G_{n}=\int_{0}^{t_{f}}dt\braket{\Psi_{n}(t)|\left[H_{\textrm{S}}(\bm{\zeta}_{0};t)-H_{0}(\bm{\zeta}_{0};t)\right]|\Psi_{0}(t)}\>.$ (29)
+Likewise, up to second order in $\mu_{\textrm{S}}$ an analogous expression for the gradient of $\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})$ reads Whi (a)
+$\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\approx-\frac{2}{\hbar^{2}}\sum_{n=1}^{\infty}\textrm{Re}\left(G_{n}\>\bm{K}_{n}^{*}\right)\>,$ (30)
+with $\bm{K}_{n}$ being an auxiliary vector function:
+$\bm{K}_{n}=\int_{0}^{t_{f}}dt\braket{\Psi_{n}(t)|\nabla_{\zeta}H_{\textrm{S}}(\bm{\zeta};t)\big.|{}_{\bm{\zeta}=\bm{\zeta}_{0}}|\Psi_{0}(t)}\>.$ (31)
+It should be stressed at this point that the evaluation of these two auxiliary functions ($G_{n}$ and $\bm{K}_{n}$) requires only the knowledge of the full Hamiltonian $H_{\textrm{S}}(\bm{\zeta};t)$ of the system and the known STA solutions for the idealized system governed by Hamiltonian $H_{0}$.
+In terms of $G_{n}$ and $\bm{K}_{n}$, the eSTA correction vector $\bm{\epsilon}$ can be expressed as
+$\bm{\epsilon}=-\frac{\left(\sum_{n=1}^{N}|G_{n}|^{2}\right)\sum_{n=1}^{N}\textrm{Re}\left(G_{n}^{\ast}\bm{K}_{n}\right)}{\left\rVert\sum_{n=1}^{N}\textrm{Re}\left(G_{n}^{\ast}\bm{K}_{n}\right)\right\rVert^{2}}\>,$ (32)
+where the infinite sum over $n$ is truncated to the first $N$ terms (i.e.
+$N$ serves as the cut-off parameter).
+From this last expression, $\bm{\epsilon}$ can straightforwardly be computed numerically provided that $G_{n}$ and $\bm{K}_{n}$ are previously calculated using Eqs.
+(29) and (31), respectively.
+A general heuristic argument can be made that provides an explanation as to why eSTA-based control protocols should be more robust (i.e.
+less sensitive) to parameter deviations than their STA-based counterparts.
+Namely, from the very assumption – inherent in the construction of the eSTA approach – that the state fidelity behaves as a quadratic function in the vicinity of its maximum [cf.
+Eq.
+(26)] and the fact that the fidelity corresponding to an eSTA scheme is higher than that of the attendant STA scheme, it follows that the derivative of the eSTA fidelity curve should be smaller than that of its STA counterpart.
+This, in turn, signifies increased robustness of the eSTA scheme.
 IV Dynamical-lattice trajectories
-In what follows, we apply the general STA results reviewed in Sec.
+In what follows, we apply the general STA- and eSTA results reviewed in Sec.
 III to coherent atom transport in the adjustable family of 2D optical lattices under consideration.
-We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice.
+We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice (see Sec.
+IV.1 below).
 We do so within the framework of a single-atom Hamiltonian with the harmonically approximated potential of Eq.
 (9).
-In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$.
-Here $\bm{\lambda}=\left(\bm{\lambda}_{\mathrm{x}},\bm{\lambda}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\lambda}_{\mathrm{x}}\equiv[\lambda^{(1)}_{\mathrm{x}},\ldots,\lambda^{(6)}_{\mathrm{x}}$] and $\bm{\lambda}_{\mathrm{y}}\equiv[\lambda^{(1)}_{\mathrm{y}},\ldots,\lambda^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively.
+We then address the same problem using the more sophisticated eSTA approach (see Sec.
+IV.2 below), which takes into account the full, anharmonic optical-lattice potential.
+In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};t)$.
+Here $\bm{\zeta}=\left(\bm{\zeta}_{\mathrm{x}},\bm{\zeta}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\zeta}_{\mathrm{x}}\equiv[\zeta^{(1)}_{\mathrm{x}},\ldots,\zeta^{(6)}_{\mathrm{x}}$] and $\bm{\zeta}_{\mathrm{y}}\equiv[\zeta^{(1)}_{\mathrm{y}},\ldots,\zeta^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively.
 The latter are assumed to satisfy the following conditions:
-$q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
-$q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{y}}\>$. (23)
-The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\lambda}_{0}\equiv\left(\bm{\lambda}_{\mathrm{x},0},\bm{\lambda}_{\mathrm{y},0}\right)^{\intercal}$, with components $\lambda^{(j)}_{\mathrm{x},0}$ and $\lambda^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$.
-Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y},0};\>t)$.
+$q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\zeta^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
+$q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\zeta^{(j)}_{\mathrm{y}}\>$. (33)
+IV.1 STA dynamical-lattice trajectories
+The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\zeta}_{0}\equiv\left(\bm{\zeta}_{\mathrm{x},0},\bm{\zeta}_{\mathrm{y},0}\right)^{\intercal}$, with components $\zeta^{(j)}_{\mathrm{x},0}$ and $\zeta^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$.
+Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y},0};\>t)$.
 We now proceed to apply the LRI-based transport theory to the approximate system Hamiltonian
@@
 the potential $V_{\mathrm{L}}^{\mathrm{x}}$, we set $\omega_{\mathrm{x}}(t)=\omega_{\mathrm{x}}$ and
-$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0,x}(t)-a_{\mathrm{x}}\right]\>.$ (26)
+$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0}^{\mathrm{x}}(t)-a_{\mathrm{x}}\right]\>.$ (36)
 Analogously, for atom transport along the $y$ direction, i.e. potential $V_{\mathrm{L}}^{\mathrm{y}}$, we have $\omega_{\mathrm{y}}(t)=\omega_{\mathrm{y}}$ and
-$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0,y}(t)\>.$ (27)
+$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0}^{\mathrm{y}}(t)\>.$ (37)
 In keeping with the standard practice Tor , we also add to $V_{\mathrm{L}}^{\mathrm{x}}$ and $V_{\mathrm{L}}^{\mathrm{y}}$ the irrelevant time-dependent global terms $m\omega^{2}_{\mathrm{x}}q^{2}_{0,x}(t)$ and $m\omega^{2}_{\mathrm{y}}q^{2}_{0,y}(t)$, respectively, which do not yield any force.
@@
 where $\mathbf{q}_{\mathrm{c}}(t)\equiv\left[q_{\mathrm{c}}^{\mathrm{x}}(t),q_{\mathrm{c}}^{\mathrm{y}}(t),0\right]^{\intercal}$ is the corresponding 2D classical-particle trajectory.
-One peculiar aspect of these last auxiliary equations is the presence of $a_{x}$ on the right-hand-side of Eq.
+One peculiar aspect of these last auxiliary equations is the presence of $a_{\mathrm{x}}$ on the RHS of Eq.
 (38), resulting from the terms linear in $x$ in the approximated potentials $V_{\mathrm{L}}$ [cf.
@@
 (9)].
-The presence of $a_{x}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor .
+The presence of $a_{\mathrm{x}}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor .
 It should be noted that Eq.
 (38) can be reduced to the form of Eq.
-(29) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{x}/\omega_{\mathrm{x}}^{2}$.
+(39) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{\mathrm{x}}/\omega_{\mathrm{x}}^{2}$.
 In this manner, the problem of atom transport in the $x$ direction, equally like the one in the $y$-direction, is reduced to a well-known type of atom-transport problems solvable via inverse engineering Tor .
-Figure 4: Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
+Figure 4: (Color online) Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
 The boundary conditions for $q_{0}^{\mathrm{x}}(t)$ and $q_{0}^{\mathrm{y}}(t)$ – along with those for for $q_{\mathrm{c}}^{\mathrm{x}}(t)$ and $q_{\mathrm{c}}^{\mathrm{y}}(t)$ consistent with Eqs.
 (38) and (39) – motivate one to seek the general solution for the path of the potential minimum in the form of the ninth-degree polynomial Hau
-$q_{0,u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (30)
+$q_{0}^{u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (40)
 where $d_{u}$ is the transport distance. From the relevant initial/boundary conditions Hau , the following results for constants $b_{n,u}$ are obtained:
@@
 $b_{7,u}$ $=$ $5040\>(t_{\mathrm{f}}\omega_{u})^{-2}+540\>,$
-$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\>.$
-The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0,x}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig.
+$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\quad(u=x,y)\>.$
+The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0}^{\mathrm{x}}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig.
 4.
+IV.2 eSTA dynamical-lattice trajectories
+Having obtained STA dynamical-lattice trajectories in the problem at hand (cf.
+Sec.
+IV.1), we now proceed to find their STA counterparts based on the methodology reviewed in Sec.
+III.3.
+The control vectors corresponding to the sought-after eSTA trajectory will be denoted in what follows by $\bm{\zeta}_{\mathrm{x,S}}$ and $\bm{\zeta}_{\mathrm{y,S}}$; accordingly, this trajectory will be described by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y,S}};\>t)$.
+The eSTA paths of the potential minimum can be expressed as
+$q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\,t)$ $=$ $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x},0};\,t)+f_{\mathrm{x}}(\bm{\alpha}_{\mathrm{x}};\,t)\>,$ (42)
+$q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y,S}};\,t)$ $=$ $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y},0};\,t)+f_{\mathrm{y}}(\bm{\alpha}_{\mathrm{y}};\,t)\>,$ (43)
+where $\bm{\alpha}:=\left(\bm{\alpha}_{\mathrm{x}},\bm{\alpha}_{\mathrm{y}}\right)^{\intercal}$ is an auxiliary control vector.
+These eSTA paths are assumed to obey the conditions $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\,jt_{f}/7)=\zeta^{(j)}_{\mathrm{x},0}+\alpha^{(j)}_{\mathrm{x}}$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{x,S}};\,jt_{f}/7)=\zeta^{(j)}_{\mathrm{y},0}+\alpha^{(j)}_{\mathrm{y}}$ ($j=1,\ldots,6$).
+The auxiliary functions $f_{u}(\bm{\alpha}_{u};\>t)$ ($u=x,y$) ought to satisfy the following boundary conditions:
+$\begin{split}&f_{u}(\bm{\alpha}_{u};0)=f_{u}(\bm{\alpha}_{u};t_{f})=0\>,\\ &f_{u}(\bm{\alpha}_{u};jt_{f}/7)=\alpha^{(j)}_{u}\quad(\>j=1,\ldots,6\>)\>,\\ &\frac{\mathrm{d^{(n)}}}{\mathrm{d}t^{(n)}}f_{u}(\bm{\alpha}_{u};t^{\prime})|_{t^{\prime}=\{0,t_{f}\}}=0\quad(\>n=1,\ldots,4\>)\>.\end{split}$ (44)
+These conditions are chosen such that $f_{u}(\bm{\alpha}_{u};\>t)$ can be controlled through $\bm{\alpha}_{u}$, satisfying at the same time the condition of continuity.
+Therefore, we choose the following Ansatz based on an eleventh-degree polynomial:
+$f_{u}(\bm{\alpha}_{u};\>t)=\sum_{n=0}^{11}\sum_{j=1}^{6}\tilde{a}^{(j)}_{u,n}\alpha^{(j)}_{u}\left(\frac{t}{t_{f}}\right)^{n}\>.$ (45)
+The concrete values of the coefficients $\tilde{a}^{(j)}_{\mathrm{x},n}$ and $\tilde{a}^{(j)}_{\mathrm{y},n}$ in these equations can be found in Table I of Hau .
+[…]
 V Calculation of atomic dynamics
-In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA-based dynamical-lattice trajectories [cf.
+In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA- and eSTA-based dynamical-lattice trajectories [cf.
 Sec.
 IV].
-This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix A).
+This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix C).
 The crucial prerequisite for computing single-atom dynamics in the problem at hand amounts to evaluating the ground-state wave-function of the relevant optical-lattice potential.
@@
 This wave function can be obtained to requisite accuracy using the plane-wave-expansion method (for details, see, e.g., Sto (b)), which entails Fourier expansions of the underlying lattice potential and the periodic part of single-particle Bloch wave functions in terms of reciprocal-lattice vectors.
-Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix B).
+Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix D).
 Given that in this study we employ the FSOM for the dynamics part, this last scheme appears as the more convenient choice.
@@
 where $\delta\dot{\mathbf{q}}_{0}(t)\equiv\dot{\mathbf{q}}_{0}(t+\delta t)-\dot{\mathbf{q}}_{0}(t)$.
-It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the RHS of the last equation, analogously to the dropped terms in Eq.
+It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the right-hand-side (RHS) of the last equation, analogously to the dropped terms in Eq.
 (47) above.
 With the aid of an analog of Eq.
-(A) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (37), switching also to the coordinate representation of the momentum operator [i.e.
+(C) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (51), switching also to the coordinate representation of the momentum operator [i.e.
 $\hat{\mathbf{p}}\rightarrow(\hbar/i)\nabla$], we arrive at the FSOM-based (second-order) time-stepping scheme for the atom-transport problem at hand:
@@
 2).
+At the same time, describing transport in the diagonal direction necessitates the use of a $45$-degree rotated square-shaped unit cell (black dashed lines in Fig.
+2).
 Yet another problem-specific feature that allows one to further reduce the level of computational difficulty in the problem at hand is closely related to its description in the comoving frame.
@@
 (50)] that depends on the acceleration $\ddot{\mathbf{q}}_{0}$ of the potential minimum.
-The spatial Fourier transformation is carried out using the FFT algorithm (cf.
-Appendix ANRc .
-To be more specific, we make use of a discrete three-dimensional $N_{x}\times N_{y}\times N_{z}$ grid with $N_{x}=200$, $N_{y}=300$, and $N_{z}=100$ points in the computational window whose dimensions (in units of the respective characteristic lenghtscales $l_{\mathrm{x}}$, $l_{\mathrm{y}}$, and $l_{\mathrm{z}}$) are $100$, 100100, and $500$, respectively.
-For the propagation of the TDSE in the comoving frame [cf.
+To carry out the spatial Fourier transformation using the FFT algorithm (cf.
+Appendix CNRc , we make use of a discrete two-dimensional $N_{\mathrm{x}}\times N_{\mathrm{y}}$ grid with $N_{\mathrm{x}}=N_{\mathrm{y}}=300$ points in the computational window whose dimensions (in units of $l_{\mathrm{x}}$ and $l_{\mathrm{y}}$) are $80$ and $80$, respectively.
+As regards the time-domain propagation of the TDSE in the comoving frame [cf.
 Eq.
-(33)], we make use of the adaptive approach NRc with a maximal relative error of $10^{-4}$; the latter could – strictly speaking – require additional (smaller in size) substeps in time.
-Yet, it turned out to be sufficient to take $N_{t}$ between $20$ and $100$ in all our runs.
+(47)], we carry it out using the adaptive approach NRc with a maximal relative error of $10^{-4}$, which could – strictly speaking – require additional (smaller) substeps in time.
+Yet, in all our runs, it proved sufficient to take $N_{t}$ between $20$ and $100$.
 VI Results and discussion
-In what follows, we quantify and discuss the efficiency of STA-based coherent single-atom transport (cf.
-Sec.
-IV above) in various optical lattices belonging to the lattice family under consideration.
-The main figure of merit used for this purpose is the transport fidelity, defined as
-$\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}\>.$ (39)
-Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix B) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec.
-V above and Appendix A).
-Figure 5: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ and potential depths using the STA scheme.
-The ratio of the lattice depths corresponds to that of a honeycomb lattice.
-Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,600,200)E_{R}$, $(400,1200,400)E_{R}$, and $(800,2400,800)E_{R}$, respectively.
-Figure 6: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme.
-The ratio of the lattice depths corresponds to that of a square lattice.
-Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
-Figure 7: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme.
-The ratio of the lattice depths corresponds to that of a linear lattice (1D chains).
-Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
-Figure 8: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme.
-The ratio of the lattice depths corresponds to that of a dimerized lattice.
-Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,400,200)E_{R}$, $(400,800,400)E_{R}$, and $(800,1600,800)E_{R}$, respectively.
-The obtained results for the transport fidelity in the optical lattices of the adjustable family under consideration are depicted in Figs.
-5-8.
-One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{f}$.
-This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$ that is proportional to the lattice depth Hau the atom being transported cannot be confined by the optical-lattice potential.
+In what follows, we quantify and discuss the efficiency of STA- and eSTA-based coherent single-atom transport (cf.
+Secs.
+IV.1 and IV.2 above, respectively) in various optical lattices belonging to the lattice family under consideration.
+The main figure of merit used for this purpose is the transport fidelity, defined as $\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}$.
+Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix D) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec.
+V above and Appendix C).
+In Sec.
+VI.1 below, we present our results for atom-transport fidelities.
+In Sec.
+VI.2 we then discuss how these results are affected in the presence of small variations in lattice depths.
+Finally, in Sec.
+VI.3 we demonstrate the superiority of our adopted transport schemes compared to approaches based on ad-hoc velocity profiles.
+VI.1 Atom-transport fidelities
+The obtained results for the transport fidelity in the optical lattices under consideration are depicted in Figs.
+512.
+One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{\mathrm{f}}$.
+This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$, which is proportional to the lattice depth Hau , the atom being transported cannot be confined by the optical-lattice potential.
 This happens as a consequence of the disappearance of the minima of the “tilted” lattice potential experienced by the atom in its comoving (non-inertial) reference frame; compared to the potential in the lab frame, the comoving-frame potential has an additional term linear in the transport coordinate Hau .
@@
 it is not special for the concrete family of lattices being considered here.
-The aforementioned breakdown of transport occurs when the maximal acceleration of an atom during transport - hereafter denoted by $|\tilde{a}_{\textrm{max}}|$ - becomes larger than $|a_{\textrm{max}}|$.
-It should be borne in mind that the actual value of $|\tilde{a}_{\textrm{max}}|$ depends on the chosen dynamical-lattice trajectory (i.e.
-the path of the potential minimum); yet, quite generally it holds that $|\tilde{a}_{\textrm{max}}|\propto t_{f}^{-2}$.
-It immediately follows from this last result that for deeper lattices – i.e., for a larger value of $|a_{\textrm{max}}|$, which is proportional to the lattice depth – the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached becomes larger than $|a_{\textrm{max}}|$ for sufficiently short total time $t_{f}$.
-In other words, for deeper lattices the breakdown of transport (i.e., the abrupt decrease of the transport fidelity) takes place for shorter times $t_{f}$.
-Our obtained numerical results for the transport fidelity [cf.
+Figure 5: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a square lattice [cf.
+Fig.
+2(a)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
+Figure 6: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=100\,l_{\mathrm{y}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a square lattice [cf.
+Fig.
+2(a)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
+The collapse of transport actually takes place when the maximal acceleration of an atom during the transport process, to be hereafter denoted by $|\tilde{a}_{\textrm{max}}|$, exceeds $|a_{\textrm{max}}|$.
+Needless to say, the actual value of $|\tilde{a}_{\textrm{max}}|$ depends upon the chosen dynamical-lattice trajectory (i.e.
+the path of the potential minimum); however, it holds quite generally that $|\tilde{a}_{\textrm{max}}|\propto\>t_{\mathrm{f}}^{-2}$.
+What follows from this last result is that for deeper lattices, which then also implies a larger value of $|a_{\textrm{max}}|$ because the latter is proportional to the lattice depth, the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached exceeds $|a_{\textrm{max}}|$ for sufficiently short transport times $t_{\mathrm{f}}$.
+In other words, for deeper lattices the collapse of transport (i.e., the precipitous decrease of the transport fidelity) takes place for shorter times $t_{\mathrm{f}}$.
+Our numerical results are consistent with this general argument, as can be verified for gradually increasing lattice depths in all figures independent on the specific transport direction or lattice configuration.
+In the opposite limit of very long transport times $t_{\mathrm{f}}$, the STA results converge to the well-known adiabatic regime.
+In this limit, essentially perfect atom transport ($\mathcal{F}\approx 1$) can be achieved for $t_{\mathrm{f}}\gg\omega_{\mathrm{u}}^{-1}$ ($u=x,y$).
+However, this behavior is only partially recovered in the eSTA approach.
+In cases where the potential minimum does not reside at the origin—corresponding to the presence of an additional linear term in the harmonic approximation, cf.
 Eq.
-(39)] are displayed in Figs.
-5-8 for honeycomb, square, linear (1D chain), and dimerized lattices, respectively, and two different transport distances ($d=100\>l_{x}$ and $d=400\>l_{x}$).
-These results are consistent with the above general features of coherent atom transport in optical lattices.
-In other words, they clearly manifest breakdown of transport at short transport times.
-On the other hand, in the opposite limit of very long transport times $t_{\mathrm{f}}$ the results coincide with the well-known adiabatic limit, i.e.
-essentially perfect atom transport ($\mathcal{F}\approx 1$) for $t_{\mathrm{f}}\gg\omega_{\mathrm{j}}^{-1}$ ($j=x,y$) is achieved.
-In particular, in the case of the honeycomb lattice, with the parameter values chosen as in Fig.
-5, the respective transport times $t_{f}$ corresponding to the onset of the breakdown of atom transport are in the range between $7\>T_{\mathrm{x}}$ and $8\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, as well as in the range between $16\>T_{\mathrm{x}}$ and $17.5\>T_{\mathrm{x}}$ for the transport distance $d=400\>l_{x}$.
-In the case of the square lattice, which corresponds to Fig.
-6, the breakdown of atom transport takes place at times in the range between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
-These results are clearly in agreement with the anticipated trend of decreasing with the increase of the lattice depth.
-For the linear (1D-chain) lattice that corresponds to Fig.
-7, the transport breakdown takes place when $t_{f}$ is between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
-Finally, in the case of the dimerized lattice with the parameter values chosen as in Fig.
-8, the onset of the transport breakdown occurs for $t_{f}$ between $10\>T_{\mathrm{x}}$ and $13\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $14\>T_{\mathrm{x}}$ and $18\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
+(12)—the adiabatic limit is reached only for sufficiently deep lattice potentials, as discussed for the honeycomb and dimer potentials at the end of this section.
+The transport fidelity corresponding to the square lattice for displacements along the $x$-direction is depicted in Fig.
+5.
+While the eSTA fidelity curve in this case exhibits a pronounced dip for shallow lattice potentials in panel (a), it nevertheless achieves a significantly improved fidelity compared to its STA counterpart for transport times $t_{\mathrm{f}}\approx 3.8\,T_{\mathrm{X}}$.
+This improvement becomes increasingly pronounced for deeper potentials, culminating in an almost uniformly high fidelity close to unity in panel (c).
+The remaining minor deviations are a shallow dip for $t_{\mathrm{f}}\approx 4.2\,T_{\mathrm{X}}$ and a slightly earlier onset of fidelity degradation for $t_{\mathrm{f}}\approx 3.2\,T_{\mathrm{X}}$.
+For even deeper lattices, these features are expected to vanish completely, yielding a consistently improved transport performance through the eSTA scheme for the square lattice.
+Owing to the high degree of symmetry of this lattice configuration [cf.
+Fig.
+2(a)], an analogous behavior is observed for displacements along the $y$-direction, as illustrated in Fig.
+6.
+For displacements along the $y$-direction in the 1D-chains lattice the corresponding STA- and eSTA fidelity curves are shown in Fig.
+8.
+What can be inferred from these results is that their qualitative behavior closely follows that of the square lattice.
+The most notable differences are the minimal achievable high-fidelity transport times, which are shifted to $t_{\mathrm{f}}\approx(6.5-7.5)\,T_{\mathrm{X}}$, and the increased lattice depth of $U_{\mathrm{\overline{X}}}$, reaching values up to $4000\,E_{R}$.
+While this depth is substantially larger than in the square-lattice case, it remains within reach of state-of-the-art experimental optical-lattice setups.
+Figure 7: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of the 1D-chains lattice [cf.
+Fig.
+2(d)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
+Figure 8: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=100\,l_{\mathrm{y}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of the 1D-chains lattice [cf.
+Fig.
+2(d)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
+By contrast to atom transport in the $y$ direction, transport along the $x$-direction of the 1D-chains lattice [for an illustration, see Fig.
+7, exhibits a qualitatively different behavior.
+Rather than displaying a sharp dip followed by an abrupt loss of fidelity, the STA transport fidelity is characterized by a more gradual degradation as the transport time $t_{\mathrm{f}}$ decreases.
+This can be attributed to the more slowly varying potential depth along the $x$-direction.
+Remarkably, the eSTA transport fidelity already provides a noticeable improvement for relatively shallow lattice depths in panel (a), albeit with a minor dip for $t_{\mathrm{f}}\approx 4.3\,T_{\mathrm{X}}$.
+For increasing lattice depths, eSTA consistently outperforms STA across the entire range of transport times, where even the previously described dip in fidelity is not visible anymore.
+It is important to notice that the sharp fidelity drop for the eSTA case observed in other geometries for small $t_{\mathrm{f}}$ is absent here; instead, eSTA transport fidelity exhibits a smoother decay as well, closely mirroring the gradual behavior of its STA counterpart.
+The STA results for transport along the $x$-direction of the honeycomb lattice, shown in Fig.
+9, similarly display a gradual reduction in transport fidelity for shorter transport times and do not exhibit a pronounced dip prior to the final drop.
+In addition, small oscillations are observed for $t_{\mathrm{f}}\approx 4.8\,T_{\mathrm{X}}$ and $t_{\mathrm{f}}\approx 5.5\,T_{\mathrm{X}}$ for all considered lattice depths, panels (a)–(c).
+These oscillations take place at times that are independent of the lattice depth, while their amplitude decreases for deeper potentials.
+These features can be attributed to the non-vanishing linear term in the harmonic approximation, Eq.
+(12), which induces minor instabilities in the eSTA-based transport scheme, particularly for shallower lattices.
+Figure 9: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a honeycomb lattice [cf.
+Fig.
+2(c)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(400,1200,400)E_{R}$, $(800,2400,800)E_{R}$, and $(1600,4800,1600)E_{R}$, respectively.
+Figure 10: (Color online) Dependence of the STA-and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=50\,l_{\mathrm{y}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a honeycomb lattice [cf.
+Fig.
+2(c)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(400,1200,400)E_{R}$, $(800,2400,800)E_{R}$, and $(1600,4800,1600)E_{R}$, respectively.
+A similar discussion applies to transport along the $y$-direction of the honeycomb lattice, illustrated by the transport-fidelity curves in Fig.
+10.
+Although the fidelity drop at short transport times occurs more abruptly and at slightly smaller values of $t_{\mathrm{f}}$ [more precisely, for $t_{\mathrm{f}}\approx(2.2-2.8)\>T_{\mathrm{X}}$], the eSTA approach nevertheless outperforms the STA scheme in this regime.
+Finally, transport in the dimer lattice along both the $x$- and $y$-directions, respectively illustrated by the fidelity curves in Figs.
+11 and 12, remains unstable even for extremely deep lattice potentials.
+This behavior can be partially attributed to the enhanced contribution of the linear term in the harmonic approximation, or, equivalently, to the large values of $x_{0}$ for this potential configuration [cf.
+Eq.
+(7)].
+These effects give rise to oscillations that are responsible for the fidelity dips observed for $t_{\mathrm{f}}\approx 6\,T_{\mathrm{X}}$ and $t_{\mathrm{f}}\approx 7.1\,T_{\mathrm{X}}$.
+In addition, the exceptionally shallow potential barrier separating the two dimer sites plays an important role.
+As illustrated in Fig.
+2(b), this barrier is extremely shallow and increases only marginally with increasing lattice depth.
+Although eSTA exhibits some improvement for deeper lattices, it does not surpass the STA performance in this case.
+We expect that eSTA would eventually outperform STA for even deeper potentials; however, computational limitations prevent us from accessing this regime even further.
+Figure 11: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a dimer lattice [cf.
+Fig.
+2(b)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(12800,25600,12800)E_{R}$, $(25600,51200,25600)E_{R}$, and $(51200,102400,51200)E_{R}$, respectively.
+Figure 12: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=50\,l_{\mathrm{y}}$ and fixed potential depths.
+The ratio of the lattice depths corresponds to that of a dimer lattice [cf.
+Fig.
+2(b)].
+Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(12800,25600,12800)E_{R}$, $(25600,51200,25600)E_{R}$, and $(51200,102400,51200)E_{R}$, respectively.
+These effects are most pronounced for eSTA-based transport along the $x$-direction, while being noticeably improved – but not entirely suppressed – for displacements along $y$, as demonstrated by the eSTA results in Fig.
+12.
+Thus, although transport in the $y$-direction exhibits enhanced stability compared to the $x$-direction, residual instabilities remain and do not vanish completely.
+By contrast, STA-based transport does not exhibit instabilities for the dimer lattice in either the $x$- or $y$-direction.
+It should be emphasized that the eSTA results for the dimer lattice require unrealistically deep potentials that are far beyond current experimental capabilities.
+Consequently, it serves primarily as an academic test case to elucidate the challenging regimes of the eSTA protocol.
+For this reason, the dimer lattice will not be considered further in the subsequent sections, which focus on experimentally relevant configurations.
+Figure 13: (Color online) Relative stability to reach an atom-transport fidelity of $0.99$, defined as $1-t_{\mathrm{real}}/t_{\mathrm{id.}}$, as a function of the instability parameter $\xi$.
+Here, $t_{0.99}$ is the transport time that first drops below $0.99$ for a given $\xi$ while $t_{\mathrm{id.}}$ is the reference time $t_{0.99}$ for $\xi=0$.
+Curves are shown for both the STA and eSTA schemes, illustrating how the transport time is affected by lattice instabilities.
+The lattice depth ratios correspond to (a) the square lattice, with $(U_{X},U_{\overline{X}},U_{Y})=(0,800,800)\,E_{R}$; (b) the 1D-chains lattice, with $(U_{X},U_{\overline{X}},U_{Y})=(100,2000,400)\,E_{R}$; (c) the honeycomb lattice $(U_{X},U_{\overline{X}},U_{Y})=(1600,4800,1600)\,E_{R}$.
+The transport distance is $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ for the square and 1D-chains lattice, and $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ for the honeycomb lattice.
+VI.2 Robustness against deviations in lattice depths
+To account for possible imperfections in envisioned experimental implementations of our proposed transport schemes, here we address the sensitivity of those schemes to deviations in lattice depths.
+It is pertinent to assume that in an actual implementation of our STA- and eSTA transport schemes the nominal potential depths $U_{j}$ ($j\in\{X,\overline{X},Y\}$) take slightly changed values $\widetilde{U}_{j}=U_{j}\left(1+\xi\right)$, where $\xi$ denotes the instability parameter.
+To simplify our analysis, we keep $\widetilde{U}_{j}$ fixed at some $\xi$-independent values and make use of those values as the actual potential depths in our evaluation of STA- and eSTA transport fidelities; we previously design STA- and eSTA dynamical-lattice trajectories (recall Sec.
+IV) with $U_{j}(\xi)\equiv\widetilde{U}_{j}/(1+\xi)$ playing the role of potential depths for each of the considered values of $\xi$.
+The advantage of this approach is that by fixing the potential depth the maximal acceleration supported by the potential remains the same for all data points (i.e.
+for different values of $\xi$), which facilitates a clearer assessment of sensitivity to deviations.
+While in the previous section a direct, one-to-one comparison between STA and eSTA schemes was central, here the focus lies on the individual stability properties of each scheme under parameter variations.
+The results of our analysis are illustrated in Fig.
+13, where the relative deviation $1-t_{0.99}/t_{\mathrm{id.}}$ is plotted as a function of the instability parameter $\xi$.
+Here, $t_{0.99}$ denotes the transport time at which the fidelity first drops below $0.99$, while the reference time $t_{\mathrm{id.}}$ corresponds to the respective scheme (STA or eSTA) evaluated at $\xi=0$.
+By construction, all curves pass through the origin.
+A notable trend in the STA results, visible across all three panels, is that increasing the instability parameter leads to a reduced transport time $t_{0.99}$, resulting in an approximately constant negative slope.
+This behavior can be attributed to the fact that decreasing $\xi$ effectively reduces the lattice depths and, consequently, the frequencies of the harmonic approximations underlying the STA construction.
+At shorter transport times $t_{\mathrm{f}}$, where higher excited states contribute more strongly, a lower effective frequency provides a better approximation of the potential far away from its minimum.
+In this sense, reducing the frequency can be interpreted as a heuristic improvement that captures the overall potential shape more faithfully than a harmonic approximation that is better suited around the origin.
+An exception to the aforementioned trend is found in the honeycomb-lattice case shown in panel (c), where a sudden increase in $t_{0.99}$ occurs around $\xi\approx-0.4$.
+This behavior likely originates from the fact that the harmonic frequencies are becoming too large to reliably represent the global structure of the underlying potential.
+Overall, the eSTA results follow a similar qualitative trend as the STA scheme, which is expected given that eSTA is constructed as a potential tailored version build on STA.
+Nevertheless, additional features emerge.
+For the square lattice in panel (a), a pronounced reduction of $t_{0.99}$ is observed near $\xi=0$, extending into the regime of negative instability parameters till around 0.2-0.2.
+In this interval, the eSTA protocol consistently outperforms the reference time $t_{\mathrm{id.}}$, before reverting to the general trend.
+Since the eSTA scheme is more strongly tailored to the reference potential, and is thus expected to be more sensitive to deviations, this region of enhanced stability does not extend over the entirety of the probed regime.
+Similar behavior is observed for the remaining lattice geometries.
+Notably, for the 1D-chains lattice shown in panel (b), the point $\xi=0$ itself lies within the region of improved stability.
+This indicates that incorporating small deviations from the reference potential in the construction of the control protocol can enhance robustness, although the extent and location of this improvement depend sensitively on the specific lattice geometry.
+VI.3 Comparison with atom transport based on ad-hoc velocity profiles
+Figure 14: (Color online) Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for fixed potential depths using the STA- and eSTA-based transport, as well as transport based on sine- and triangular (ad-hoc) velocity profiles.
+The ratio of the lattice depths corresponds to that (a) the square lattice with $(U_{X},U_{\overline{X}},U_{Y})=(0,800,800)E_{R}$; (b) the 1D-chains lattice with $(U_{X},U_{\overline{X}},U_{Y})=(200,4000,800)E_{R}$; (c) the honeycomb lattice with $(U_{X},U_{\overline{X}},U_{Y})=(1600,4800,1600)E_{R}$.
+The transport distance is $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ for the square and 1D-chains lattice, and $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ for the honeycomb lattice.
+In the following, we conclude our analysis of STA and eSTA-based atom transport by comparing the results for the atom-transport fidelity obtained using these approaches with those corresponding to ad-hoc (independent of the optical-lattice potential) velocity profiles.
+More specifically yet, we make comparisons between our STA- and eSTA results for atom transport to those based on the sine-shaped and triangular velocity profiles.
+The respective time-dependent paths $q_{0}^{\mathrm{triang}}(t)$ and $q_{0}^{\mathrm{sin}}(t)$ of the potential minimum corresponding to the triangular and sine-shaped velocity profiles are given by
+$q_{0}^{\mathrm{triang}}(t)$ $=\begin{cases}v_{0}\,t^{2}/t_{f}\>,&\;\;0\leq t\leq t_{f}/2\\ v_{0}\left(2\,t-t_{f}/2-t^{2}/t_{f}\right)\>,&\;\;t_{f}/2<t\leq t_{f}\end{cases}\>,$ (53)
+$q_{0}^{\mathrm{sine}}(t)$ $=\frac{v_{0}}{2}\left[t-\frac{\sin\left(2\pi\frac{t}{t_{f}}\right)}{2\pi}\>t_{f}\right]\>,$ (54)
+where $v_{0}=2d/t_{f}$ is the maximal velocity during transport over the distance $d$.
+The atom-transport approach based on the triangular velocity profile [cf.
+Eq.
+(53)] is often referred to as the bang-bang approach Tor ; Che (b); Din .
+This approach is accompanied by significant motional-heating effects in the fast-transport regime, a phenomenon inextricably linked with the presence of discontinuities in its corresponding acceleration profile.
+Consequently, such an approach leads to relatively inefficient atom transport in most practically relevant cases Hic .
+Somewhat better results than the ones originating from the bang-bang approach are obtained for atom transport based on the sine-shaped velocity profile [cf.
+Eq.
+(54)]; this is intimately related to the fact that its attendant acceleration of an atom remains continuous throughout the transport process.
+By contrast to atom-transport approaches based on ad-hoc velocity profiles, the inverse-engineering based STA approach, and its eSTA counterpart to an even larger extent, make use of the specific form of the optical-lattice potential in question to obtain a tailored dynamical-lattice trajectory.
+For this reason, it is plausible to expect that these last two approaches should lead to much more time-efficient and robust atom transport than those based on pre-selected velocity profiles.
+For the optical-lattice family under consideration, a comparison between transport based on ad-hoc velocity profiles and transport based on STA- and eSTA approaches is illustrated in Fig.
+14 for the cases of the square-, 1D-chains-, and honeycomb lattice.
+As is evident from Fig.
+14, the obtained results for the transport fidelity in the case of the triangular velocity show strong oscillations.
+The results obtained in the case of the sine-shaped velocity profile are closer – but still inferior – to those originating from the STA- and eSTA approaches.
+From the results obtained using the latter approaches it can be inferred that only one major drop in fidelity takes place in the case of STA-based transport in the square lattice [cf.
+Fig.
+14(a)], while for the remaining two lattice configurations [cf.
+Fig.
+14(b) and (c)] such drops are completely absent for both STA and eSTA transport fidelities.
 VII Summary and Conclusions
-To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices.
-We first specified concrete arrangements of acousto-optic modulators that give rise to the moving-lattice effect in these systems in an arbitrary direction.
-We subsequently designed the corresponding dynamical-lattice trajectories using shortcuts to adiabaticity (STA).
-Through numerical propagation of the relevant time-dependent Schrödinger equation – using the previously computed STA-based dynamical-lattice trajectories – we calculated the efficiency ofhe resulting coherent atom transport as described by the atom-transport fidelity.
-We demonstrated that the STA approach enables fast atom transport in the considered family of optical lattices, for a wide range of lattice depths and transport directions.
-In addition, we analyzed the implications of the atom-transport processes for neutral-atom quantum computing based on collisional two-qubit entangling gates.
+To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices which includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases.
+We first specified concrete arrangements of acousto-optic modulators that give rise to the dynamical(moving)-lattice effect in these systems in an arbitrary direction.
+We subsequently designed the corresponding dynamical-lattice trajectories using both shortcuts to adiabaticity (STA) in the form of inverse engineering based on a Lewis-Riesenfeld type dynamical invariant, as well as their enhanced version (eSTA).
+Through the numerical propagation of the relevant time-dependent Schrödinger equation – using the previously obtained STA and eSTA dynamical-lattice trajectories – we quantified the efficiency of the resulting coherent atom transport as described by the atom-transport fidelity.
+By computing transport fidelities for wide ranges of values of relevant parameters (lattice depths, transport distances, etc.) and different transport directions for four different lattice configurations, we showed that – with the notable exception of the dimerized lattice – the eSTA method consistently outperforms its STA counterpart; this superiority of the eSTA approach is manifested both in terms of the achievable transport times and the robustness of the resulting transport against small variations of optical-lattice depths.
+For the sake of completeness, we also demonstrated that our adopted eSTA approach to atom transport is far superior to alternative approaches to modeling atom transport in optical lattices, more precisely the approaches based on ad-hoc velocity profiles.
 By contrast to previous theoretical investigations of coherent atom transport, based mostly on simplified scenarios, in the present work we addressed this phenomenon based on the full, anharmonic optical-lattice potential.
@@
 Furthermore, our study also complements well the recent works on quantum-state control of Bose-Einstein condensates in optical lattices Amr .
-Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems, the first one of which has been undertaken quite recently Hwa .
+Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems (including those based on the eSTA approach), the first one of which has been undertaken quite recently Hwa .
 Acknowledgements.
@@
 This research was supported by the Deutsche Forschungsgemeinschaft (DFG) – SFB 1119 – 236615297.
-Appendix A FSOM-based numerical propagation of TDSEs
+Appendix A Derivation of the expression for $G_{n}$
+In the following, we provide the derivation of an expression that serves as a prerequisite for the calculation of the scalar auxiliary function $G_{\mathbf{n}}$ [cf.
+Eq.
+(29)] in the problem under consideration.
+For brevity, the two-dimensional index $\mathbf{n}\equiv(n_{\mathrm{x}},n_{\mathrm{y}})$ will hereafter be used whenever possible.
+Furthermore, the evaluated integrals can be calculated with respect to a specific coordinate, however, they can readily be transferred to their corresponding counterparts.
+The excitation number of the integral indicates the coordinate over which the integration was performed.
+The transport modes corresponding to the approximated Hamiltonian $H_{\mathrm{L},0}(t)$ [cf.
+Eq.
+(34)], which describes transport within the xyx-y plane, have the form characteristic of a 2D harmonic oscillator.
+In the coordinate representation, those transport modes are given by a 2D generalization of the expression in Eq.
+(22), in which the Lewis-Riesenfeld phase is given by Eq.
+(23).
+In other words, we have
+$\begin{split}\braket{\bm{r}|\Psi_{\mathbf{n}}}&=\exp\left[-\frac{i}{\hbar}\left(E_{\mathbf{n}}t+\frac{m}{2}\int_{0}^{t}\mathrm{d}t^{\prime}\left[\dot{q}_{\mathrm{c,x}}(t^{\prime})^{2}+\dot{q}_{\mathrm{c,y}}(t^{\prime})^{2}\right]\right)\right]\exp\left(im\left[\dot{q}_{\mathrm{c,x}}(t)x+\dot{q}_{\mathrm{c,y}}(t)y\right]/\hbar\right)\\ &\hskip 14.22636pt\times\left(2^{n}n_{\mathrm{x}}!n_{\mathrm{y}}!\pi^{2}\right)^{-1/2}\mathrm{H}_{n_{\mathrm{x}}}\left(\frac{x-q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}\sqrt{2}}\right)\mathrm{H}_{n_{\mathrm{y}}}\left(\frac{y-q_{\mathrm{c}}^{\mathrm{y}}(t)}{l_{\mathrm{y}}\sqrt{2}}\right)\\ &\hskip 14.22636pt\times\left(2l_{\mathrm{x}}\,l_{\mathrm{y}}\right)^{-1/2}\,\exp\left(-\frac{[x-q_{\mathrm{c}}^{\mathrm{x}}(t)]^{2}}{4l_{\mathrm{x}}^{2}}\right)\,\exp\left(-\frac{[y-q_{\mathrm{c}}^{\mathrm{y}}(t)]^{2}}{4l_{\mathrm{y}}^{2}}\right)\>.\end{split}$ (55)
+Here $E_{\mathbf{n}}\equiv\hbar\omega_{\mathrm{x}}(n_{\mathrm{x}}+1/2)+\hbar\omega_{\mathrm{y}}(n_{\mathrm{y}}+1/2)-ma_{\mathrm{x}}^{2}/(2\omega_{\mathrm{x}}^{2})$ is the $n$-th energy eigenvalue, which contain a constant-energy offset due to the presence of a linear term in $x$ in the approximated potential of Eq.
+(9); $l_{\mathrm{j}}\equiv\sqrt{\hbar/\left(2m\omega_{j}\right)}$ is the harmonic-oscillator zero-point length along the direction $j$ ($j=x,y$), while $H_{n}(x)$ ($n\in\mathbb{N}$) is the Hermite polynomial Cho .
+By inserting the expression from Eq. (55) into Eq. (29), we obtain the following expression for the scalar auxiliary function:
+$\begin{split}G_{\mathbf{n}}=&\int_{0}^{t_{\mathrm{f}}}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}Y\int_{-\infty}^{\infty}\mathrm{d}X\,\frac{\exp\left[i\left(\omega_{\mathrm{x}}n_{\mathrm{x}}+\omega_{\mathrm{y}}n_{\mathrm{y}}\right)t\right]}{\sqrt{2^{n}n_{\mathrm{x}}!n_{\mathrm{y}}!\pi^{2}}}\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\mathrm{H}_{n_{\mathrm{y}}}\left[Y_{\mathrm{C}}(t)\right]\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\exp\left[-Y_{\mathrm{C}}^{2}(t)\right]\\ &\times\Bigg\{U_{\mathrm{L}}\left[\sqrt{2}\,X_{0}(t)\,l_{\mathrm{x}},\sqrt{2}\,Y_{0}(t)\,l_{\mathrm{y}}\right]-\frac{\hbar}{2}\left[\omega_{\mathrm{x}}X_{0}^{2}(t)+\omega_{\mathrm{y}}Y_{0}^{2}(t)\right]-\sqrt{2}\,m\,a_{\mathrm{x}}\,l_{\mathrm{x}}\,X_{0}(t)-V_{\mathrm{d,0}}\Bigg\}\>.\end{split}$ (56)
+In the last equation, we introduced the new, dimensionless variables $X=x/(l_{\mathrm{x}}\sqrt{2})$ and $Y=y/(l_{\mathrm{y}}\sqrt{2})$, as well as the new functions $X_{0}(t):=X-q_{0}^{\mathrm{x}}(t)/(l_{\mathrm{x}}\sqrt{2})$, $Y_{0}(t):=Y-q_{0}^{\mathrm{y}}(t)/(l_{\mathrm{y}}\sqrt{2})$, $X_{\mathrm{C}}(t):=X-q_{\mathrm{c}}^{\mathrm{x}}(t)/(l_{\mathrm{x}}\sqrt{2})$ and $Y_{\mathrm{C}}(t):=Y-q_{\mathrm{c}}^{\mathrm{y}}(t)/(l_{\mathrm{y}}\sqrt{2})$.
+The evaluation of the integral in Eq.
+(56) reduces to the calculation of several auxiliary intermediate integrals.
+The first class of integrals that have to be considered are those with integrands that do not have additional dependence on $X_{0}$; they are given by
+$\begin{split}W_{n_{\mathrm{x}}}(t)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\\ &=\int_{-\infty}^{\infty}\mathrm{d}X\sum_{m=0}^{\infty}\frac{1}{m!}\,\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{m}\sum_{l=0}^{n_{\mathrm{x}}}\begin{pmatrix}n_{\mathrm{x}}\\ l\end{pmatrix}\mathrm{H}_{l}\left(X\right)\left[-\sqrt{2}\>\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}\right]^{n_{\mathrm{x}}-l}\mathrm{H}_{m}\left(X\right)\exp\left(-X^{2}\right)\,\\ &=\sqrt{\pi}\,\delta_{n_{\mathrm{x}},0}\>.\end{split}$ (57)
+This integral was evaluated by making use of the Faá di Bruno representation of Hermite polynomials Wei (b)
+$\mathrm{H}_{m}(x)=(-1)^{m}\sum_{k_{1}+2k_{2}=m}\frac{m!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}\left(2\mathnormal{x}\right)^{k_{1}}\>.$ (58)
+Integrals of the form of Eq.
+(57) vanish based on the well-known orthogonality condition for Hermite polynomials for all but the zero-excitation mode.
+This results in constant offsets only contributing in the trivial case of $\mathbf{n}=(0,0)$, however, for the calculation of $G_{\mathbf{n}}$ this case does not occur [cf.
+Eq.
+(29)].
+In a similar way, we can obtain the final result for an integrand that has a term linear in $X_{0}$. This quantity can be straightforwardly derived using the following expression:
+$\begin{split}L_{n_{\mathrm{x}}}(t)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\,X_{0}(t)\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\\ &=\int_{-\infty}^{\infty}\mathrm{d}X\sum_{m=0}^{\infty}\frac{1}{m!}\,\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{m}\sum_{l=0}^{n_{\mathrm{x}}}\begin{pmatrix}n_{\mathrm{x}}\\ l\end{pmatrix}\mathrm{H}_{l}\left(X\right)\left[-\sqrt{2}\>\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}\right]^{n_{\mathrm{x}}-l}\\ &\hskip 14.22636pt\times\Bigg[\frac{1}{2}\,\mathrm{H}_{m+1}\left(X\right)-\frac{q_{0}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}\,\mathrm{H}_{m}\left(X\right)+m\mathrm{H}_{m-1}\left(X\right)\Bigg]\exp\left(-X^{2}\right)\,\\ &=\sqrt{\pi}\,\left[\sqrt{2}\,\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}\right]^{n_{\mathrm{x}}-1}\sum_{l=0}^{n_{\mathrm{x}}}\begin{pmatrix}n_{\mathrm{x}}\\ l\end{pmatrix}(-1)^{n_{\mathrm{x}}-l}\left\{l+\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]}{2\,l_{\mathrm{x}}^{2}}\right\}\>.\end{split}$ (59)
+This result is obtained using the same steps as in the evaluation of the integral in Eq. (57) above, with an additional usage on the identity Cho
+$x\,\mathrm{H}_{m}(x)=\frac{1}{2}\>\mathrm{H}_{m+1}(x)+m\,\mathrm{H}_{m-1}(x)\>.$ (60)
+It is straightforward to prove inductively that the only non-vanishing contributions are those with $n_{\mathrm{x}}\in\{0,1\}$. Therefore, the final result is given by
+$L_{n_{\mathrm{x}}}(t)=\sqrt{\pi}\left[\delta_{n_{\mathrm{x}},1}+\delta_{n_{\mathrm{x}},0}\,\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)}{\sqrt{2}\,l_{\mathrm{x}}}\right]\>.$ (61)
+The integration involving the $X_{0}^{2}(t)$-dependent terms of Eq. (56) entails no new methods of integration and can readily be obtained through the following
+$\begin{split}Q_{n_{\mathrm{x}}}(t)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\,X_{0}^{2}(t)\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\\ &=\int_{-\infty}^{\infty}\mathrm{d}X\sum_{m=0}^{\infty}\frac{\mathrm{H}_{m}\left(X\right)}{m!}\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\,l_{\mathrm{x}}}\right]^{m}\exp\left(-X^{2}\right)\,\\ &\hskip 14.22636pt\times\left[X^{2}-\sqrt{2}\>X\frac{q_{0}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}+\frac{{q_{0}^{\mathrm{x}}}^{2}(t)}{2\,l_{\mathrm{x}}^{2}}\right]\,\sum_{l=0}^{n_{\mathrm{x}}}\begin{pmatrix}n_{\mathrm{x}}\\ l\end{pmatrix}\mathrm{H}_{l}\left(X\right)\left[-\sqrt{2}\>\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}\right]^{n_{\mathrm{x}}-l}\\ &=\sqrt{\pi}\,\left[\sqrt{2}\,\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}\right]^{n_{\mathrm{x}}}\,\\ &\hskip 14.22636pt\times\sum_{l=0}^{n_{\mathrm{x}}}\begin{pmatrix}n_{\mathrm{x}}\\ l\end{pmatrix}(-1)^{n_{\mathrm{x}}-l}\Bigg\{\frac{1}{4}\,l\,(l-1)\,\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}\right]^{-2}-l\,q_{0}^{\mathrm{x}}(t)\,{q_{\mathrm{c}}^{\mathrm{x}}(t)}^{-1}+l\Bigg\}\>,\end{split}$ (62)
+where use has been made of the binomial theorem. The above integral evaluates to
+$\begin{split}Q_{n_{\mathbf{x}}}(t)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\,X_{0}^{2}(t)\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\\ &=\sqrt{\pi}\,\left[\delta_{n_{\mathrm{x}},0}\,\frac{{q_{0}^{\mathrm{x}}}^{2}(t)}{2\>l_{\mathrm{x}}^{2}}+\delta_{n_{\mathrm{x}},1}\,\frac{\sqrt{2}\,q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)}{l_{\mathrm{x}}}+2\>\delta_{n_{\mathrm{x}},2}\right]\>.\end{split}$ (63)
+The second complex class of integrals that we need to evaluate is formed by dependents through a trigonometric relation on $X_{0}(t)$.
+The first integral of this class that we consider is the following
+$C_{n_{\mathrm{x}}}^{a}\left(t\right)=\int_{-\infty}^{\infty}\mathrm{d}X\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\cos\left[\sqrt{2}\,a\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,X_{0}(t)\right]\>,$ (64)
+with $a\in\mathbb{R}$.
+This integral can be evaluated in the same manner as the previously considered ones, beginning by rewriting the cosine term in terms of exponential functions.
+In this way, one obtains the expression
+$C_{n_{\mathrm{x}}}^{a}\left(t\right)$ $=$ $\frac{(-1)^{n_{\mathrm{x}}}}{2}\,\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}2^{k_{1}}\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\int_{-\infty}^{\infty}\mathrm{d}X\,X^{l}\,\exp\left[-\left(X-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)^{2}\right]$ (65)
+$\times$ $\Bigg\{\exp\left[i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,\left(X-\frac{q_{0}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)\right]+\exp\left[-i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,\left(X-\frac{q_{0}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)\right]\Bigg\}\>.$
+The integral in Eq. (65) can be computed by ordering the different terms and employing the following general result:
+$\int_{-\infty}^{\infty}\mathrm{d}x\,x^{n}\,\exp\left(-\mathrm{a}x^{2}+\mathrm{b}x+\mathrm{c}\right)=\exp\left(\frac{b^{2}}{4a}+c\right)\sum_{k=0}^{\lfloor n/2\rfloor}\begin{pmatrix}n\\ 2k\end{pmatrix}\left(\mathrm{\frac{b}{2a}}\right)^{n-2k}\frac{\Gamma\left(k+1/2\right)}{\mathrm{a}^{k+1/2}}\>,$ (66)
+where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}\>e^{-t}\mathrm{d}t$ is the gamma function (hereafter used only for a real-valued argument xxCho .
+This leads to the final expression for the cosine-dependent term
+$\begin{split}C_{n_{\mathrm{x}}}^{a}\left(t\right)&=\frac{(-1)^{n_{\mathrm{x}}}}{2}\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}\,2^{k_{1}}\,\exp\left(-\frac{a^{2}k_{\mathrm{L}}^{2}l_{\mathrm{x}}^{2}}{2}\right)\,\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\\ &\hskip 14.22636pt\times\sum_{\lambda=0}^{\lfloor l/2\rfloor}\begin{pmatrix}l\\ 2\lambda\end{pmatrix}\Gamma\left(\lambda+1/2\right)\,\Bigg\{\exp\left(ia\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\right)\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}+ia\,\frac{k_{\mathrm{L}}l_{\mathrm{x}}}{\sqrt{2}}\right]^{l-2\lambda}\\ &\hskip 14.22636pt+\exp\left(-ia\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\right)\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}-ia\,\frac{k_{\mathrm{L}}l_{\mathrm{x}}}{\sqrt{2}}\right]^{l-2\lambda}\Bigg\}\>.\end{split}$ (67)
+Similarly, to evaluate the sine-function counterpart of the integral in Eq. (65), one follows analogous steps. In this manner, we arrive at the following final result:
+$\begin{split}S_{n_{\mathrm{x}}}^{a}\left(t\right)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\sin\left[\sqrt{2}\,a\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,X_{0}(t)\right]\\ &=(-1)^{n_{\mathrm{x}}-1}\frac{i}{2}\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}\,2^{k_{1}}\,\exp\left(-\frac{a^{2}k_{\mathrm{L}}^{2}l_{\mathrm{x}}^{2}}{2}\right)\,\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\\ &\hskip 14.22636pt\times\sum_{\lambda=0}^{\lfloor l/2\rfloor}\begin{pmatrix}l\\ 2\lambda\end{pmatrix}\Gamma\left(\lambda+1/2\right)\,\Bigg\{\exp\left(ia\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\right)\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}+ia\,\frac{k_{\mathrm{L}}l_{\mathrm{x}}}{\sqrt{2}}\right]^{l-2\lambda}\\ &\hskip 14.22636pt-\exp\left(-ia\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\right)\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}\>l_{\mathrm{x}}}-ia\,\frac{k_{\mathrm{L}}l_{\mathrm{x}}}{\sqrt{2}}\right]^{l-2\lambda}\Bigg\}\>.\end{split}$ (68)
+It should be noticed that the only differences of this last result compared to Eq. (67) are the change of sign in the curly brackets and the additional prefactor of (i)(-i).
+Another term that has to be evaluated is similar to Eq. (64), but instead of a simple cosine term, it contains a cosine-squared, term. This integral is given by
+$\widetilde{C}_{n_{\mathrm{x}}}\left(t\right)$ $=$ $\int_{-\infty}^{\infty}\mathrm{d}X\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\,\exp\left[-X^{2}_{\mathrm{C}}(t)\right]\cos^{2}\left[\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,X_{0}(t)\right]$ (69)
+$=$ $\frac{1}{4}\,\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}2^{k_{1}}\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\int_{-\infty}^{\infty}\mathrm{d}X\,X^{l}\,\exp\left[-\left(X-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)^{2}\right]$
+$\times$ $\Bigg\{\exp\left[i\,2^{3/2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,\left(X-\frac{q_{0}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)\right]+\exp\left[-i\,2^{3/2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,\left(X-\frac{q_{0}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right)\right]+2\Bigg\}\>.$
+Since no new techniques were used to evaluate this integral, we just state the final result, which is given by
+$\begin{split}\widetilde{C}_{n_{\mathrm{x}}}\left(t\right)&=\frac{(-1)^{n_{\mathrm{x}}}}{4}\,\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}2^{k_{1}}\,\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\,\sum_{\lambda=0}^{\lfloor l/2\rfloor}\begin{pmatrix}l\\ 2\lambda\end{pmatrix}\Gamma\left(\lambda+1/2\right)\,\exp\left(-2\,k_{\mathrm{L}}^{2}\,l_{\mathrm{x}}^{2}\right)\\ &\times\Bigg\{\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}+i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\right]^{l-2\lambda}\exp\Big\{+i\,2\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\Big\}\\ &+\hskip 11.38092pt\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}-i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\right]^{l-2\lambda}\exp\Big\{-i\,2\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\Big\}\Bigg\}+\frac{\sqrt{\pi}}{2}\,\delta_{n_{\mathrm{x}},0}\>.\end{split}$ (70)
+The same holds true for the sine-counterpart of the above integral, which results in an almost identical expression:
+$\begin{split}\widetilde{S}_{n_{\mathrm{x}}}\left(t\right)&=\int_{-\infty}^{\infty}\mathrm{d}X\,\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\,\exp\left[-X_{\mathrm{C}}(t)^{2}\right]\sin^{2}\left[\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\,X_{0}(t)\right]\\ &=\frac{(-1)^{n_{\mathrm{x}}+1}}{4}\,\sum_{k_{1}+2k_{2}=n_{\mathrm{x}}}\frac{n_{\mathrm{x}}!}{k_{1}!k_{2}!}(-1)^{k_{1}+k_{2}}2^{k_{1}}\,\sum_{l=0}^{k_{1}}\begin{pmatrix}k_{1}\\ l\end{pmatrix}\left[-\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}\right]^{k_{1}-l}\,\sum_{\lambda=0}^{\lfloor l/2\rfloor}\begin{pmatrix}l\\ 2\lambda\end{pmatrix}\Gamma\left(\lambda+1/2\right)\,\exp\left(-2\,k_{\mathrm{L}}^{2}\,l_{\mathrm{x}}^{2}\right)\\ &\times\Bigg\{\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}+i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\right]^{l-2\lambda}\exp\Big\{+i\,2\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\Big\}\\ &+\hskip 11.38092pt\left[\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}-i\,\sqrt{2}\,k_{\mathrm{L}}\,l_{\mathrm{x}}\right]^{l-2\lambda}\exp\Big\{-i\,2\,k_{\mathrm{L}}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]\Big\}\Bigg\}+\frac{\sqrt{\pi}}{2}\,\delta_{n_{\mathrm{x}},0}\>.\end{split}$ (71)
+By combining the obtained results for the integrals discussed so far and making use of the orthogonality of Hermite polynomials, the scalar auxiliary function $G_{\mathbf{n}}$ can be reduced to the time integral
+$\begin{split}G_{\mathbf{n}}=&-\int_{0}^{t_{\mathrm{f}}}\mathrm{d}t\,\frac{\exp\left[i\left(\omega_{\mathrm{x}}n_{\mathrm{x}}+\omega_{\mathrm{y}}n_{\mathrm{y}}\right)t\right]}{\sqrt{2^{n}n_{\mathrm{x}}!n_{\mathrm{y}}!\pi}}\Bigg\{\eta_{0,\bm{n}}(t)+\eta_{1,\bm{n}}(t)+\eta_{2,\bm{n}}(t)\\ &+\hbar\sqrt{\pi}\Bigg[\omega_{\mathrm{x}}\delta_{n_{\mathrm{y}},0}\,\left(\delta_{n_{\mathrm{x}},1}\,\frac{q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)}{\sqrt{2}l_{\mathrm{x}}}+\delta_{n_{\mathrm{x}},2}\right)+\omega_{\mathrm{y}}\delta_{n_{\mathrm{x}},0}\,\left(\delta_{n_{\mathrm{y}},1}\,\frac{q_{\mathrm{c}}^{\mathrm{y}}(t)-q_{0}^{\mathrm{y}}(t)}{\sqrt{2}l_{\mathrm{y}}}+\delta_{n_{\mathrm{y}},2}\right)\Bigg]\\ &+\sqrt{2\,\pi}\,m\,a_{\mathrm{x}}\,l_{\mathrm{x}}\,\delta_{n_{\mathrm{x}},1}\,\delta_{n_{\mathrm{y}},0}\Bigg\}\>,\end{split}$ (72)
+where $\eta_{i,\bm{n}}(t)$, $i\in\left\{0,1,2\right\}$ stand for the following functions of time:
+$\begin{split}\eta_{0,\bm{n}}(t)=&\>U_{X}\,\widetilde{S}_{n_{\mathrm{x}}}(t)\,\delta_{n_{\mathrm{y}},0}+U_{Y}\,\widetilde{C}_{n_{\mathrm{y}}}(t)\,\delta_{n_{\mathrm{x}},0}\>,\\ \eta_{1,\bm{n}}(t)=&\>2\gamma\>\sqrt{\frac{U_{X}\,U_{Y}}{\pi}}\>S_{n_{\mathrm{x}}}^{a=1}(t)\,C_{n_{\mathrm{y}}}^{a=1}(t)\>\cos\phi\>,\\ \eta_{2,\bm{n}}(t)=&\>\frac{U_{\overline{X}}}{2}\left[C_{n_{\mathrm{x}}}^{a=2}(t)\>\cos\theta-S_{n_{\mathrm{x}}}^{a=2}(t)\>\sin\theta\right]\,\delta_{n_{\mathrm{y}},0}\>.\end{split}$ (73)
+The integration over time in Eq. (72), the final step in the evaluation of $G_{\mathbf{n}}$, has to be carried out numerically.
+Appendix B Derivation of the expression for $\mathbf{K}_{\mathbf{n}}$
+In the following, we derive an expression that can be used as the basis for a numerical evaluation of the second (vector) auxiliary function $\mathbf{K}_{\mathbf{n}}$ [cf.
+Eq.
+(31)] in the problem at hand.
+To derive the desired expression for $\mathbf{K}_{\mathbf{n}}$, the gradient with respect to $\bm{\zeta}$, and subsequently to $\bm{\alpha}$, of the Hamiltonian $H_{\mathrm{DW}}$ is needed.
+Since the differentiations are straightforward, we just state the resulting equation for $\mathbf{K}_{\mathbf{n}}$:
+$\begin{split}\mathbf{K}_{\mathbf{n}}&=\,\int_{0}^{t_{\mathrm{f}}}\,\mathrm{d}t\,\,\int_{-\infty}^{\infty}\,\mathrm{d}Y\,\int_{-\infty}^{\infty}\,\mathrm{d}X\frac{1}{\sqrt{2^{n}n_{\mathrm{x}}!n_{\mathrm{y}}!\pi^{2}}}\exp\left[i\left(n_{\mathrm{x}}\omega_{\mathrm{x}}+n_{\mathrm{y}}\omega_{\mathrm{y}}\right)t\right]\\ &\hskip 14.22636pt\times\mathrm{H}_{n_{\mathrm{x}}}\left[X_{\mathrm{C}}(t)\right]\mathrm{H}_{n_{\mathrm{y}}}\left[Y_{\mathrm{C}}(t)\right]\,\exp\left[-X_{\mathrm{C}}^{2}(t)\right]\exp\left[-Y_{\mathrm{C}}^{2}(t)\right]\\ &\hskip 14.22636pt\times\bm{\nabla}_{\bm{\alpha}}\underbrace{\left\{U_{X}[\sqrt{2}\>X_{0}(t)l_{\mathrm{x}}]+U_{Y}[\sqrt{2}\>Y_{0}(t)l_{\mathrm{y}}]+U_{XY}[\sqrt{2}\>X_{0}(t)l_{\mathrm{x}},\sqrt{2}\>Y_{0}(t)l_{\mathrm{y}}]\right\}}_{U_{\mathrm{D}}}\Big|_{\bm{\alpha}=\bm{0}}\>.\end{split}$ (74)
+The derivations of the potential contributions are given by
+$\bm{\nabla}_{\bm{\alpha}}U_{X}[x-q_{0}^{\mathrm{x}}(\alpha)]$ $=\>\>\>U_{X}k_{\mathrm{L}}\sin\left[2k_{\mathrm{L}}(x-q_{0}^{\mathrm{x}})\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}-U_{\overline{X}}k_{\mathrm{L}}\sin\left[2k_{\mathrm{L}}(x-q_{0}^{\mathrm{x}})+\theta\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}\>,$
+$\bm{\nabla}_{\bm{\alpha}}U_{Y}[y-q_{0}^{\mathrm{y}}(\alpha)]$ $=-U_{Y}k_{\mathrm{L}}\sin\left[2k_{\mathrm{L}}(y-q_{0}^{\mathrm{y}})\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{y}}\>,$
+$\bm{\nabla}_{\bm{\alpha}}U_{XY}[x-q_{0}^{\mathrm{x}}(\alpha),y-q_{0}^{\mathrm{y}}(\alpha)]$ $=-2\,k_{\mathrm{L}}\,\gamma\,\cos\phi\,\sqrt{U_{X}U_{Y}}\Big\{\sin\left[k_{\mathrm{L}}(x-q_{0}^{\mathrm{x}})\right]\,\sin\left[k_{\mathrm{L}}(y-q_{0}^{\mathrm{y}})\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{y}}$
+$\hskip 99.58464pt-\cos\left[k_{\mathrm{L}}(x-q_{0}^{\mathrm{x}})\right]\,\cos\left[k_{\mathrm{L}}(y-q_{0}^{\mathrm{y}})\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}\Big\}\>.$ (75)
+In fact, all of the needed integrals have already been carried out in the evaluation of the first auxiliary function (cf. Appendix A). Thus, here we just state the final result
+$\begin{split}\mathbf{K}_{\mathbf{n}}&=\int_{0}^{t_{\mathrm{f}}}\,\mathrm{d}t\,\frac{\exp\left[i\left(n_{\mathrm{x}}\omega_{\mathrm{x}}+n_{\mathrm{y}}\omega_{\mathrm{y}}\right)t\right]}{\sqrt{2^{n}n_{\mathrm{x}}!n_{\mathrm{y}}!\pi}}\Big[\tilde{\bm{\eta}}_{0,\bm{n}}(t)\,\delta_{n_{\mathrm{y}},0}+\tilde{\bm{\eta}}_{1,\bm{n}}(t)\,\delta_{n_{\mathrm{x}},0}+\tilde{\bm{\eta}}_{2,\bm{n}}(t)\Big]\>,\end{split}$ (76)
+where $\tilde{\bm{\eta}}_{0,\bm{n}}(t)$, $\tilde{\bm{\eta}}_{1,\bm{n}}(t)$ and $\tilde{\bm{\eta}}_{2,\bm{n}}(t)$ are vector functions of time given by:
+$\begin{split}\tilde{\bm{\eta}}_{0,\bm{n}}(t)&=\>\>\>U_{X}\,k_{\mathrm{L}}\,S_{n_{\mathrm{x}}}^{a=2}(t)\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}-U_{\overline{X}}\,k_{\mathrm{L}}\,\left[S_{n_{\mathrm{x}}}^{a=2}(t)\cos\theta+C_{n_{\mathrm{x}}}^{a=2}(t)\sin\theta\right]\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}\>,\\ \tilde{\bm{\eta}}_{1,\bm{n}}(t)&=-U_{Y}\,k_{\mathrm{L}}\,S_{n_{\mathrm{y}}}^{a=2}(t)\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{y}}\>,\\ \tilde{\bm{\eta}}_{2,\bm{n}}(t)&=-2\,k_{\mathrm{L}}\,\gamma\,\cos\phi\,\sqrt{\frac{U_{X}U_{Y}}{\pi}}\Big[S_{n_{\mathrm{x}}}^{a=1}(t)S_{n_{\mathrm{y}}}^{a=1}(t)\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{y}}-C_{n_{\mathrm{x}}}^{a=1}(t)C_{n_{\mathrm{y}}}^{a=1}(t)\bm{\nabla}_{\bm{\alpha}}f_{\mathrm{x}}\Big]\>.\end{split}$ (77)
+The remaining integral over time can only be evaluated numerically, as has already been done in the evaluation of the first auxiliary function $G_{\mathbf{n}}$ [cf. Eq. (III.3)] in Appendix A.
+Appendix C FSOM-based numerical propagation of TDSEs
 In the following, we briefly review the essential ingredients of the use of the FSOM for solving TDSEs.
@@
 If carried out using the fast Fourier transform (FFT) algorithm NRc , an elementary step in these transformations entails $\mathcal{O}(N_{s}\log_{2}N_{s})$ operations.
-Appendix B Ground-state calculation via the ITE approach
+Appendix D Ground-state calculation via the ITE approach
 Here we provide a short introduction into the ITE approach, which established itself as one of the most robust computational methods for obtaining ground states of quantum systems And ; Fei .
-Its relationship to the FSOMa - its counterpart utilized for computing dynamics - bears close analogy to the connection between Lanczos-type diagonalizations and Chebyshev-propagator methods Stojanović and Salom (2019).
+Its relationship to the FSOMa - its counterpart utilized for computing dynamics - bears close analogy to the connection between Lanczos-type diagonalizations Sto (3) and Chebyshev-propagator methods Stojanovic+Salom:19 .
 It is prudent to start our review of the ITE approach using the abstract, representation-independent notation for quantum states.
@@
 The remaining steps in the application of the ITE approach within this representation are carried out in complete analogy with the above representation-independent expressions (84) and (85).
+Appendix E Derivation of dimensionless units used in the numerical simulation
+To facilitate an accurate numerical simulation of atom transport, it is pertinent to introduce dimensionless units derived from the intrinsic parameters of the system.
+This new choice of units leads to improved numerical stability and avoids potential underflow- or overflow issues that may arise when working directly with physical units.
+In particular, we define characteristic time-, length-, and energy scales based on the harmonic reference frequency of the system, which are then used consistently throughout the simulation.
+More specifically yet, we set:
+$t_{\mathrm{sys}}\equiv\omega_{\mathrm{x}}^{-1}\>,\hskip 14.22636ptl_{\mathrm{sys}}\equiv\sqrt{2}\>l_{\mathrm{x}}\>,\hskip 14.22636ptE_{\mathrm{sys}}\equiv\hbar\omega_{\mathrm{x}}\>.$
+In particular, the advantage of including the additional factor of $\sqrt{2}$ in the reference length of the approximate Hamiltonian is that it implicitly rescales the kinetic term such that the effective mass is set to unity, rather than $1/2$; this choice eliminates unnecessary prefactors in the Hamiltonian and leads to a simpler and more transparent numerical implementation.
+With the above definitions in place, the energy of the system can be expressed as
+$U_{j}=\tilde{U}_{j}E_{\mathrm{sys}}\equiv U_{j}^{R}E_{\mathrm{R}}\qquad(\>j\in\{X,Y,\overline{X}\}\>)\>.$ (87)
+In our simulations, we specify the energy scales in terms of the recoil energy $E_{\mathrm{R}}$ [cf.
+Sec.
+II.1] by setting $U_{j}^{R}$ externally, while the numerical implementation relies on the corresponding dimensionless harmonic reference units $\tilde{U}_{X}$.
+However, being defined through the wavenumber $k_{\mathrm{L}}$, the recoil energy depends explicitly on the characteristic length scale of the system, resulting in a nontrivial connection between the energy scales.
+By expressing the wavenumber $k_{\mathrm{L}}$ in system units as $k_{\mathrm{L}}=\tilde{k}/l_{\mathrm{sys}}$, the recoil energy becomes directly related to the dimensionless wavenumber via
+$\frac{\hbar^{2}k_{\mathrm{L}}^{2}}{2m}=\frac{\hbar^{2}\tilde{k}^{2}}{2ml_{\mathrm{sys}}^{2}}\equiv\frac{\tilde{k}^{2}}{2}\>\hbar\omega_{\mathrm{x}}\>.$ (88)
+Further, by making use of the reference timescale $t_{\mathrm{sys}}=\omega_{\mathrm{x}}^{-1}$, together with the derived harmonic frequency of the approximated potential [cf. Eq. (11)], we obtain
+$\omega_{\mathrm{x}}^{2}$ $=\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\>g_{\mathrm{x}}$
+$=2\tilde{k}^{2}\tilde{U}_{X}g_{\mathrm{x}}\frac{\hbar\omega_{\mathrm{x}}}{l_{sys}^{2}m}$ (89)
+$=2\tilde{k}^{2}\tilde{U}_{X}g_{\mathrm{x}}\>\omega_{\mathrm{x}}^{2}\>,$
+where we define the geometric factor in the $x$-direction as
+$g_{\mathrm{x}}\equiv\left(1-\frac{U_{\overline{X}}}{U_{X}}\right)\cos^{2}(k_{\mathrm{L}}x_{0})\>.$ (90)
+It is worthwhile noting that in Eq.
+(E) above $\omega_{x}^{2}$ enters multiplicatively on both sides, thus implying that the prefactor on the RHS must equal unity.
+This is a direct consequence of defining the system variables in terms of the effective frequency corresponding to the harmonically-approximated optical potential.
+As a result, one obtains a simple relation between the wavenumber and the potential depth in system units:
+$\tilde{k}=\frac{1}{\sqrt{2\tilde{U}_{X}g_{\mathrm{x}}}}\>.$ (91)
+Along with the previously derived relation for the recoil energy in system units [cf. Eq. (88)], we obtain
+$\tilde{U}_{X}=U_{X}^{R}\>\frac{E_{R}}{E_{\mathrm{sys}}}\equiv U_{X}^{R}\>\frac{\hbar^{2}\tilde{k}^{2}}{2m\>\hbar\omega_{\mathrm{x}}l_{\mathrm{sys}}^{2}}\>,$ (92)
+which further leads to
+$\tilde{U}_{X}=U_{X}^{R}\>\frac{\tilde{k}^{2}}{2}=U_{X}^{R}\>\frac{1}{4\tilde{U}_{X}g_{\mathrm{x}}}\>.$ (93)
+By solving the last equation for $\tilde{U}_{X}$, we arrive at the sought-after relation between the potential depth in terms of system and recoil energies in the $x$ direction:
+$\tilde{U}_{X}=\sqrt{\frac{U_{X}^{R}}{4g_{\mathrm{x}}}}.$ (94)
+Analogous relations hold for the two remaining potential depths, i.e.
+$\tilde{U}_{Y}=\sqrt{\frac{U_{Y}^{R}\kappa_{Y}}{4g_{\mathrm{x}}}}\>,\quad\tilde{U}_{\overline{X}}=\sqrt{\frac{U_{\overline{X}}^{R}\kappa_{\overline{X}}}{4g_{\mathrm{x}}}}\>,$ (95)
+where, in order to further simplify these last equations, we introduced the relative potential depths:
+$\kappa_{Y}=\frac{U_{Y}}{U_{X}}\>,\hskip 14.22636pt\kappa_{\overline{X}}=\frac{U_{\overline{X}}}{U_{X}}\>.$ (96)
+It is important to notice that the parameters in Eq. (95) still depend on the geometric factor $g_{\mathrm{x}}$ and not on their respective geometric factors that could be derived analogously.
+To summarize, the above derivation constitutes a self-consistent mapping between experimentally relevant recoil-energy scales and the dimensionless parameters used in the numerical simulation.
+By specifying the recoil energies $U_{j}^{R}$, the corresponding potential depths in system units $\tilde{U}_{j}$ are uniquely determined, while simultaneously fixing the dimensionless wavenumber $\tilde{k}$ through the harmonic reference frequency.